### v1

# Mahyar Habibi,1 Dirk Hovy,2 Carlo Schwarz1

Abstract
There is an ongoing debate about how to moderate toxic speech on social media and how content moderation affects online discourse. We propose and validate a methodology for measuring the content-moderation-induced distortions in online discourse using text embeddings from computational linguistics. We test our measure on a representative dataset of 5 million US political Tweets and find that removing toxic Tweets distorts online content. This finding is consistent across different embedding models, toxicity metrics, and samples. Importantly, we demonstrate that content-moderation-induced distortions are not caused by the toxic language. Instead, we show that, as a side effect, content moderation shifts the mean and variance of the embedding space, distorting the topic composition of online content. Finally, we propose an alternative approach to content moderation that uses generative Large Language Models to rephrase toxic Tweets to preserve their salvageable content rather than removing them entirely. We demonstrate that this rephrasing strategy reduces toxicity while minimizing distortions in online content.
Keywords: social media, content moderation, content distortions, toxicity, embeddings.
Significance Statement
Automated content moderation on social media platforms aims to reduce toxic content but is frequently criticized for distorting online discourse. However, the extent of these distortions is unknown. We present a novel measure to quantify content-moderation-induced distortions in online discourse. Analyzing 5 million US political tweets, we show that content moderation causes significant changes in the conceptual space of online debate, comparable to removing entire topics (e.g., abortion, BLM) from the platform. Finally, we show that simply rephrasing toxic content with large language models can reduce toxicity while minimizing distortions in online discourse. Our findings offer critical insights and concrete strategies for policymakers and platforms to balance content distortion with harm reduction, highlighting the trade-offs of content moderation.
1 Introduction
The widespread proliferation of hateful and inflammatory content online has become an increasing concern for users, policymakers, and online platforms. As growing evidence shows that hateful online content can lead to real-life violence (Müller and Schwarz, 2021, 2022b; Bursztyn et al., 2019; Du, 2023; Cao et al., 2023), platforms have increasingly resorted to content moderation efforts to stem the tide of hateful content online. Prominent examples include the removal of Facebook accounts associated with the far-right group Proud Boys in October 2018 (e.g., NBC-News, 2018), the deletion of Alex Jones’ Twitter account in the aftermath of the Sandy Hook shooting (e.g., BBC, 2018), or most prominently, the suspension of Donald Trump’s Twitter account after the attack on the US capitol on January 6th, 2021 (e.g., Twitter, 2021; NYT, 2021). [^1] Trump’s account was only reinstated after the takeover of Twitter by Elon Musk (The Guardian, 2022). As part of the staff cuts at Twitter, Elon Musk also fired most of the content moderators on Twitter (e.g., USA-Today, 2022) Over the years, lawmakers have also started to introduce regulation of online platforms that demand the removal of Toxic online content. For example, Germany’s “Netzwerksdurchsetzungsgesetz” (BBC, 2017), the UK’s “Online Safety Bill”(e.g, Reuters, 2023b), and the “Digital Services Act” of the EU (e.g., Reuters, 2023a) mandate that online platforms are responsible for the content that is circulating on them and therefore have to take content moderation measures. As of 2020, laws that mandate removing toxic content from social media platforms had been passed in at least 25 countries (Justitia, 2020).
On the one hand, concerns about hateful content and the increased demand for content moderation have motivated extensive research on automated hate speech detection (e.g., Waseem and Hovy, 2016; Hanu and Unitary team, 2020; Hartvigsen et al., 2022; Bianchi et al., 2022) and the effectiveness of content moderation efforts (e.g., Chandrasekharan et al., 2017; Jhaver et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022; Jiménez Durán et al., 2022; Müller and Schwarz, 2022a). One upshot of this research is that algorithms are susceptible to false positives, often triggered by swear words or otherwise innocent words that happen to often be found in the context of hate speech (e.g., Attanasio et al., 2022). On the other hand, the expansions of content moderation have been criticized as restrictions to free speech and as a distortion to online discourse (e.g., Tworek, 2021; Eidelman and Ruane, 2021; United Nations Human Rights, 2018). In particular, potentially biased applications of content rules have attracted growing criticism from politicians (e.g., Samples, 2019; Vogels et al., 2020; The Texas Tribune, 2022). Further, content moderation might also falsely target people who share encounters with racism (Lee et al., 2024). As a result, online platforms face a dilemma of seemingly contradictory objectives, which they need to balance in their content moderation efforts. This dilemma of balancing the removal of inflammatory content and the preservation of plurality of opinions is exacerbated by the fact that extensive disagreements exist about how these two objectives should be weighted.
It is worth highlighting that this trade-off would persist even if the “ground truth” of hate speech was perfectly known, i.e., if there was an unbiased and universally agreed-upon measure of hate speech. [^2] In a recent paper, Solomon et al. (2024) argue that US partisan differences over what constitutes hate speech might be overstated. Even in this hypothetical scenario, content moderation would distort online content if specific topics and issues were more frequently discussed using toxic language. The trade-off only vanishes in the highly unlikely case in which the toxicity of online discourse is entirely unrelated to its content.
The dilemma of content moderation is further complicated by a lack of measures to quantify the distortions to online content. While many methods exist to identify hateful content (see examples above), to the best of our knowledge, no measures exist to quantify the effect of content moderation-induced changes in online content. In this paper, we narrow this gap by putting forward and validating a measure for the distortions of online content. We formalize the idea of content distortions based on changes in the statistical moments (mean and variance) of the semantic space.
As a first step, we approximate the semantic space using text embeddings based on Transformer models (Vaswani et al., 2017). These embeddings represent texts as projections into a high-dimensional Euclidean space, where their semantic similarity to other texts determines their position. Texts with similar meanings will have embeddings that are closer together than semantically unrelated texts. Embeddings have become the de-facto standard text representation in natural language processing (NLP). As of 2024, the original Vaswani et al. (2017) has received over 110,000 citations, and Transformer models have proven highly successful in countless applications (e.g., Zhu et al., 2020; Strudel et al., 2021; Han et al., 2021; Radford et al., 2023), as they have been shown to capture text semantics better than previous approaches. Transformer models also form the basis of large language models and modern machine translation. Text embeddings (as opposed to count-based methods) also have proven highly successful in computational social science (e.g., Ash and Hansen, 2023; Garg et al., 2018; Kozlowski et al., 2019; Card et al., 2022).
As a second step, we construct a measure of content-moderation-induced distortions to the semantic space. Our measure is based on the Bhattacharyya distance, a widely used metric for the similarity of probability distributions. The Bhattacharyya distance captures distortions in the semantic space based on shifts in the mean and the variance of the multivariate normal distributed embedding vectors. A great advantage of this approach is that it is content-agnostic, i.e., it does not involve any choice of which content is worth preserving. [^3] Note that our argument is not that all content is worth preserving, but rather that our measure does not involve any explicit choice in this regard. Further, the measure is scalable to large datasets and entire platform ecosystems.
We validate the measure’s potential on a representative sample of 5 Million US political Tweets. Our analysis proceeds in four steps. First, we show that removing toxic Tweets leads to measurable distortions in online content. Importantly, no such distortions occur if Tweets are removed at random. This result persists independent of the embedding model, toxicity score, or Twitter sample. To the best of our knowledge, our paper is the first to provide a measure to quantify this trade-off empirically. We can also show that toxicity-based content moderation distorts both the mean and the variance of the semantic space. Additionally, we find that the variance of the semantic space is reduced by increased content moderation.
Second, we show that the content moderation-induced distortions are not driven by the position of toxic language in the embedding space. To this end, we build on the literature on the debiasing of text embeddings (e.g., Bolukbasi et al., 2016), and create projections of the embeddings space that are orthogonal to the toxicity scores. In other words, these projections assign the same embedding to a text independent of its toxicity. We find that our results remain virtually identical with these modified embeddings. This finding suggests that it is not the toxic language itself but rather that some topics that are disproportionally discussed in a toxic manner.
Third, we shed more light on the content moderation-induced distortions of online content using a Top2Vec topic model (Angelov, 2020). This allows us to characterize which topics are disproportionally affected by content moderation. Further, to benchmark our findings, we show that removing all Tweets with a Toxicity above a threshold of 0.8 produces distortions in online content comparable to removing 4 out of 67 topics from the data at random.
Lastly, we propose an alternative approach to content moderation that, instead of removing toxic content, rephrases Tweets in a way that strips them of their toxic language while leaving the core message intact. For this analysis, we use the language generation capabilities of large language models (LLMs) (e.g., Argyle et al., 2023). Using the rephrased Tweets, we can show that this alternative form of content moderation can measurably reduce toxicity while mitigating distortions to the semantic space.
Our results break new ground by proposing a measure of content distortion that allows us to quantify the inherent trade-off in content moderation. Given the fundamental importance of this trade-off, our measure is of immediate policy relevance when it comes to the moderation of toxic and fake news content. The economic theory of multitask models (Holmstrom and Milgrom, 1991; Feltham and Xie, 1994) predicts that if principles have to choose between different objectives, only one of which is measurable, the effort will be focused on the measurable tasks. In other words, in the absence of readily available measures of content distortions, online platforms, and lawmakers are likely to put far greater emphasis on removing hateful content at the “cost” of distorting online content. Hence, our measure represents an essential puzzle piece in the debate on content moderation.
2 Data and Methods
2.1 Representative US Twitter Data
For our main analysis, we use the Tweets from a representative sample of US Twitter (Siegel et al., 2021). [^4] Among others, these data were also used in Müller and Schwarz (2022a); Bose et al. (2024). The sample was created by querying the Twitter user accounts API for random numbers between 1 and $2^{32}$, the largest possible Twitter user ID at the time of collection. [^5] Twitter later switched user IDs to 64-bit. If the API returned a user account associated with the random number, the authors confirmed that the user was located in the United States. We collected the Tweets of 432,882 out of 498,901 users whose accounts were still active at the beginning of 2022. In total, this yields a dataset of ca. 400 Million Tweets. For our analysis, we removed non-English Tweets and Retweets, including those containing only links. [^6] The filtering of non-English Tweets is based on the language tags from the Perspectives API. We then finetune a BERTweet model (Nguyen et al., 2020) for the classification of political Tweets (see Appendix A for details). We use the tuned model to classify all Tweets as either political or apolitical.
We create two samples for our analysis. Our main analysis is based on a sample of 5 million randomly-drawn political Tweets. [^7] As we show in our analysis, removing Tweets at random has no impact on our measure. The results therefore would be identical if we used all Tweets instead of the subsample. However, a sample of 5 Million is computationally more efficient to handle. Secondly, for robustness checks, we create a sample of 1 million randomly-drawn Tweets independent of whether they are political or apolitical. Together, the two samples provide a good approximation of political and overall Twitter content in the United States. In robustness checks we additionally consider a sample of 1 Million German and Italian Tweets from Jiménez Durán et al. (2022); Lupo et al. (2024).
For our content moderation analysis, we assign three toxicity scores to each Tweet based on different models. Our baseline model is Google’s Perspectives API (Per, ). The Perspectives API has become one of the standard tools for toxicity analysis and is used by several platforms for content moderation (e.g., NYT, 2016; Forbes, 2019; Delgado, 2019). For each Tweet, the API returns six scores measuring different toxicity dimensions (toxicity, severe toxicity, identity attack, insult, profanity, and threat). The toxicity scores range from 0 (non-toxic) to 1 (highly toxic) and are roughly interpretable as the share of users who would judge a message as toxic. As is standard in the literature, we focus on Toxicity scores. [^8] As we show in our robustness tests, the results are very similar if we use any of the other toxicity dimensions. We create similar toxicity scores using the Detoxify package (Hanu and Unitary team, 2020) and the Moderation API from OpenAI (OpenAI, 2024). We show that both of these alternative models produce toxicity scores that are strongly positively correlated with our baseline measure (see Appendix Figure A.2). Additional details on the data and the toxicity measures can be found in Appendix A. Appendix Table A.3 provides examples of toxic Tweets in our data (Warning: The examples contain offensive language.)
2.2 Measuring Distortions to Online Content
As described in the introduction, our measure of content distortions is built around the idea of the semantic space. To build intuition, imagine that the semantic content of a text can be represented as a vector in potentially infinite-dimensional semantic space. In this space, texts that talk about the same issue or hold the same opinion are close together, while texts about other issues or diverging opinions are far apart. Our measure of content distortions will measure to what extent the removal of toxic content changes the mean and the variance of the semantic space. The measure will be small if the semantic space after the removal of toxic content is very similar to the original semantic space prior to content moderation. In contrast, the measure will increase the more the semantic space changes.
Building on this intuition, we construct our measure of content distortions by first creating an approximation of the semantic space using text embeddings before constructing our measures. We describe each of these steps in the following. First, we create embeddings for each of the Tweets in our data using the BERTweet model (Nguyen et al., 2020). [^9] We also provide robustness checks for embeddings created by the RoBerta and DeBerta models. The BERTweet model, trained on a large English-language Twitter corpus, transforms the text of each Tweet into a 768-dimensional vector. In this way, the BERTweet model generates an approximation of the semantic space. [^10] Additional details on the BERTweet model can be found in Section B.1.
These types of embedding vectors are crucial for countless NLP tasks, such as text classification, similarity calculation, summarization, translation, generation, and question-answering (e.g., Devlin et al., 2018; Radford et al., 2019; Lewis et al., 2019). In line with the above intuition, the individual dimensions of the embedding vector capture semantic differences between Tweets, i.e., those closer in the embedding space are more similar. At the end of this step, we are left with a $N\times D$ matrix $\mathbf{X}$, where $N$ is the number of Tweets, and $D$ is the number of embedding dimensions.
As the second step, we construct a measure of content distortion based on the embeddings matrix $\mathbf{X}$. Our measure is based on the Bhattacharyya distance (BCD), a commonly used metric for comparing probability distributions. As the embeddings are approximately normally distributed (see Appendix Figure B.1), we use the formula for Bhattacharyya distance for multivariate normal distributions (e.g., Abou-Moustafa and Ferrie, 2012). For the case of two multivariate normal distributions $\mathcal{N}_{1}(\mu_{1},\Sigma_{1})$ and $\mathcal{N}_{2}(\mu_{2},\Sigma_{2})$, the BCD is defined as:
$BCD(\mathcal{N}_{1},\mathcal{N}_{2})=\frac{1}{8}(\mu_{1}-\mu_{2})^{T}\Sigma^{-1}(\mu_{1}-\mu_{2})+\frac{1}{2}\left(\frac{\det\Sigma}{\sqrt{\det\Sigma_{1}\cdot\det\Sigma_{2}}}\right)$ (1)
where $\Sigma=\frac{\Sigma_{1}+\Sigma_{2}}{2}$. The Bhattacharyya intuitively consists of two additive terms. The first term is the squared Mahalanobis distance (Mahalanobis, 1936). It measures the inverse-variance-weighted distortions to the mean of the multivariate normal distribution, while the second term measures distortions to the variance. The $\det\Sigma$ is the so-called generalized variance index (GVI) (Wilks, 1932), which provides a multivariate extension of the standard statistical variance measure based on the mean squared deviation.
To calculate the BCD, we calculate both the means ($\mu$) and the variance-covariance matrix ($\Sigma$) of the embedding matrix $\mathbf{X}$ [^11] As the calculation of the dot product $XX^{\prime}$ is computationally unfeasible, we instead use a maximum likelihood estimator of the empirical covariance matrix using the procedure implemented in scikit learn. In tests on a subsample of the data, we confirmed that the approximate is very close to the analytical solution of the variance-covariance matrix. The diagonal elements of the variance-covariance matrix capture the dispersion of Tweets along the embedding dimensions. Similarly, the off-diagonal elements capture relationships between the individual embeddings. [^12] The individual entries in $\Sigma(\mathbf{X})$ are relatively small, which can lead to integer underflow when calculating the GVI. To avoid this issue, we multiply the embedding matrix $\mathbf{X}$ by 100 before calculating the BCD. Given that this equally affects all components of the BCD, this has no bearing on our results.
In this way, BCD provides a unidimensional measure that summarizes the overall distortions of the embeddings matrix $\mathbf{X}$. The BCD has several properties that make it a desirable measure for our application:
  1. 1.
    The BCD has an intuitive interpretation as both the mean and the variance capture key parameters of the multivariate normal distribution.
  2. 2.
    The BCD does not require any choice regarding which content is more valuable than others. It solely depends on the initial embedding space. In this sense, the BCD is content agnostic.
  3. 3.
    Another key advantage of the BCD is its computational tractability in high-dimensional settings. The BCD allows us to approximate the distortions of the embedding space in a computationally feasible manner, a crucial aspect given the expansive nature of online platforms and our data. [^13] Many other alternative methods, such as the ones that involve the computation of convex hulls, become computationally infeasible in high dimensions settings.
3 Results
In the following, we use the BCD to establish four sets of results. First, we analyze how the content is distorted when we remove highly toxic Tweets from the data. This analysis allows us to simulate the effect of more stringent content moderation. Second, we demonstrate that the content moderation-induced distortions are not driven by the toxic language of the Tweets. Third, we investigate what underlies the changes in the BCD. Fourth, we propose an alternative approach to content moderation that reduces toxicity while preserving the content of Tweets.
3.1 Removal of Toxic Content and Distortions of Online Discourse
We simulate the effects of more stringent content moderation by removing toxic Tweets based on varying toxicity thresholds from our data. We then analyze if removing toxic Tweets leads to distortions of online content as measured by the BCD. We will compare these content-moderation-induced changes in the semantic space to changes in the BCD if we instead remove the same number of Tweets from the data at random. Note that this analysis intentionally abstracts from any form of behavioral reaction on the side of the users, as we are interested in the direct effects of content moderation. This analysis should provide a realistic content moderation benchmark as the Perspectives API is used in real-world applications, and several studies of content moderation have used toxicity thresholds to delineate toxic content (e.g., Gehman et al., 2020; Han and Tsvetkov, 2020; Rieder and Skop, 2021; Hede et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022).
The results from this analysis are presented in Figure 1. The x-axis indicates the toxicity threshold above which we remove Tweets. We also report the share of Tweets that remain in the sample in parentheses below. The y-axis shows our measure of the distortion of the semantic space after the removal of toxic content relative to the original semantic space as measured by the BCD. It is immediately apparent that the removal of toxic content distortions to the semantic space (orange line). The findings indicate that distortions become more severe the lower we set the toxicity threshold. Importantly, removing Tweets from the data at random does not impact the BCD (blue line). This highlights that the increase in the BCD is not a mechanical consequence of a smaller sample of Tweets but rather driven by the changing composition of online discourse due to content moderation. We interpret this result as evidence that removing toxic distorts online content as measured by the embedding space.
Figure 1: Content Distortions and Removal of Toxic Content
Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis. The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random. The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
To benchmark the resulting distortions, we compare them to an approximation of the maximum possible distortion one could create by removing a specific number of Tweets. For this exercise, we remove the Tweets with the largest Euclidean distance from the centroid of the semantic space. Appendix Figure 1 shows the results from this exercise. We find that the removal of toxic content reaches around 20% of the maximum possible distortion. In Section 3.2, we provide additional benchmarks of the distortions relative to the direct removal of topics.
Robustness
We conduct several robustness exercises for this finding. First, we reproduce our finding using alternative embeddings based on the widely-used RoBerta and DeBerta models (see Figure C.2). Second, we base content moderation on the alternative Toxicity dimensions from the Perspectives API, the toxicity scores from the Detoxify classifiers (Hanu and Unitary team, 2020), and the Moderation API from OpenAI (see Figure C.3). Lastly, we repeat our analysis using samples of 1 million: 1) general Tweets (without filtering for political content), 2) German Tweets, and 3) Italian Tweets (see Figure C.4).
We discuss all of these robustness checks in greater detail in Appendix C. To summarize, none of these changes make any qualitative difference to our findings. Independent of the embedding model, the toxicity scores, or the sample, removing toxic content reduces the plurality of online discourse. Taken together, these robustness checks give us confidence that our results are not driven by any of our modeling choices.
3.2 What Drives Distortions to Online Content
We next aim to shed light on the factors that drive the content-moderation-induced distortions of the semantic space we have documented. First, we investigate the individual components of the BCD and show that both the mean as well as the variance of the embedding space get distorted by content moderation. Second, we show that the increase in the BCD is not driven by the use of toxic language itself. Third, we provide further evidence for the content-moderation-induced distortions of the topics of online debate.
Decomposing the Bhatacharya Distance
As previously discussed, the BCD consists of two additive components (see eq. 1), the first of which captures distortions to the mean, while the latter captures distortions to the variance. We analyze to what extent each of these components contributes to the previously documented increases in the BCD. The results from this analysis are shown in Panel (a) of Figure 2. We observe that while both the first and the second component contribute to the increase in the BCD, the shift in the variance explains around 96% of the increase in the BCD, while the shift in the mean only explains 4%.
Figure 2: Decompostion of Bhatacharya Distance
(a) Mean vs Variance Shift
(b) Change in GVI
Notes: Panel (a) shows a composition of the BCD into its two additive components. The blue line plots the first component (mean shifts) and the orange line plots the second component (variance shifts). Panel (b) shows changes in the generalized variance index (GVI) after excluding toxic and random Tweets from the sample. We report the log ratio of the new relative to the old GVI.
Another interesting finding we can establish is that content moderation reduces the variance of the embedding space as measured by GVI ($\det\Sigma$) (see Panel (b) Figure 2). This subfigure plots the natural logarithm of the ratio of the GVI after content moderation relative to the original GVI. It is immediately apparent that the GVI decreases significantly with the removal of toxic content (orange line). Similar to our previous results, no change in the GVI occurs if content is removed at random (blue line). As the GVI captures the dispersion of content, the GVI is small if all Tweets are very similar and, therefore, close to each other in the semantic space. In contrast, if Tweets are widely dispersed in the semantic space, the variance will be large. The results suggest that content moderation appears to remove outliers in the semantic space. Removing such “outlier” content is arguably costly for the plurality of online speech.
Accounting for the Toxicity Dimension of the Embedding Space
The content moderation-induced changes in the semantic space could be driven by two factors. On the one hand, the toxic language itself could occupy a particular part of the semantic space. Removing toxic content would then mechanically increase the BCD, but this would not necessarily indicate a “real” distortions to online content. On the other hand, toxic Tweets might discuss specific issues and topics that are underrepresented online, independently of their toxicity.
We investigate this question by removing the toxicity component from the embedding space. To do so, we build on the literature on the debiasing of embeddings (e.g., Bolukbasi et al., 2016; Liang et al., 2020) and create orthogonal projections of the embeddings matrix $\mathbf{X}$ with respect to the Toxicity scores. Our baseline approach uses linear regressions of the following form:
$x_{d}=\alpha+\mathbf{Tox^{\prime}\beta}+\epsilon_{d}$
where $x_{d}\in\mathbf{X}$ is one of the $D$ embedding dimensions. $\mathbf{Tox}$ is a matrix containing 1000 indicators for the permilles of the toxicity distribution. [^14] We chose a non-parametric transformation of the toxicity scores to flexibly account for any potential non-linearities. As we show in a robustness check, the findings are almost identical if we instead residualize linearly with regard to the toxicity score. We estimate these regressions for all embedding dimensions $d\in D$ and replace the embedding dimensions with the regression residuals. This leaves us with a residualized matrix $\mathbf{\tilde{X}}$, which is orthogonal to the toxicity scores. In other words, we remove any variation in the embeddings of Tweets that can be explained by their toxicity. Any remaining variances in the embeddings should, therefore, capture differences in Tweets’ content independent of their toxicity.
We then repeat our previous analysis based on the residualized embedding matrix $\mathbf{\tilde{X}}$. Figure 3 visualizes the two different approaches to account for a Tweet’s toxicity. More specifically, we use regressions and either residualize the embeddings using permilles or the linear toxicity score.
We find that removing the toxicity component from the embedding space hardly makes any difference to our findings. The overall patterns are very close to our original results. The BCD sharply increases once toxic Tweets are removed from the data. This suggests that the observed changes in the BCD do not appear to be driven by the toxicity of the Tweets but rather by their content.
Figure 3: Controlling for Toxicity
Notes: The figure shows the BCD, derived from toxicity-debiased Tweet embeddings, after the exclusion of toxic and random Tweets from the sample. Linear and Percentile Residualization adjust each dimension of Tweets embeddings by using the residuals of the regression of the embedding values against the toxicity score and toxicity permilles of Tweets.
Content-Moderation-Induced Topic Shifts
As the last piece of evidence on the content-moderation-induced shifts in online content, we analyze how the topic composition of Tweets is shifting with increased content moderation. Additionally, we provide evidence that, in line with our original motivation, the removal of topics is associated with increases in the BCD. The latter result also allows us to benchmark the content-moderation-induced increases in the BCD relative to the direct removal of topics from the embedding space.
In Panel (a) of Figure 4, we show the topics that would be affected most by content moderation. Perhaps unsurprisingly, we find that Tweets containing insults against political figures are most affected by content moderation. [^15] Note that even highly uncivil insults against public figures represent a form of opinion expression. Moreover, we also find that Tweets discussing the topics of “Racism”, “Trump”, or the “Black Lives Matter (BLM)” movement are heavily affected by content moderation, thereby distorting the topic composition of online debate. [^16] The finding on the removal of discussions of racism aligns well with the work of Lee et al. (2024). It is also important to highlight that our results do not hinge on shifts in any one topic, i.e., the results are very virtually identical if we do not consider Tweets containing insults (see Appendix Figure C.5). Further, we can show that content moderation distorts the semantic space even within topics (unreported).
Figure 4: Content Moderation and Topic Shifts
(a) Topics Shifts
(b) BCD and Removal of Topics
Notes: The figure visualizes the effect of content moderation on changes in topics as generated by the Top2Vec algorithm. Panel (a) shows which topics are most heavily affected by content moderation. Panel (b) shows that removing 1, 2, 3, 4, or 5 topics from the data at random leads to increases in the BCD. The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
In Panel (b), we lastly show that removing topics leads to increasing changes in the BCD. For this figure, we calculate the BCD for the cases in which we remove 1, 2, 3, 4, or 5 topics from the data at random. For this exercise, we implement 25 random draws of the indicated number of topics. We then remove Tweets from the selected topics from the data and calculate the BCD. The figure then reports the average BCD resulting from the 25 draws. We also report the average share of Tweets that remain in the data after the removal of topics in brackets on the x-axis.
We find that the BCD increases progressively with the removal of additional topics. This exercise also allows us to benchmark the content-moderation-induced distortions we documented in Figure 1. We find that removing all Tweets with a toxicity score above 0.8 from the data is akin to removing four topics from the data at random. Both of these interventions result in a BCD of ca. 0.7. It is also worth highlighting that removing four deletes twice as many Tweets from the data (6%) as content moderation with a threshold of 0.8 (3%). This suggests that the per-Tweet distortions of the semantic space are more severe when removing content based on toxicity.
Taken stock, the previous exercises suggest that the reduction in BCD are not driven by the toxic language itself, but rather the fact that removing Tweets with toxic language as a byproduct 1) shifts the mean of the embeddings space, 2) reduces the variance of the semantic space, and 3) removes topics from the online debate. In the last part of the paper, we propose an alternative approach to mitigate some of the content-moderation-induced distortions.
3.3 An Alternative Approach to Content Moderation
We propose an alternative approach to tackle the issue of online toxicity that has the potential to mitigate some of the distortions in online discourse. Instead of removing toxic content outright, content moderators could instead use the language generation capabilities of large language models and rephrase the message of Tweets using less toxic language. This makes it possible to reduce the toxicity of online content while keeping the original content as far as possible intact. [^17] The detoxification of online content has recently also become a task tackled by computer scientists in the TextDetox 2024 competition.
Note that we do not argue that all content should necessarily remain on the platform. There might well be content and thoughts that should not be admitted to be shared on online platforms. Our argument rather is that there exist significant amounts of content that violate the content standards of social media platforms (e.g., grave insults) but, at the same time, might contain political expression or information that might be worth preserving. In such cases, rephrasing might be a suitable approach. For example, the rephrasing of messages has been shown to facilitate communication in partisan politics (Argyle et al., 2023).
We test the potential of such an approach using GPT4o-mini to rephrase Tweets with a toxicity score above 0.5. We provide additional details on the rephrasing prompt in Section B.2. We also show examples of rephrased Tweets in Table A.3. Overall, GPT4o-mini can rephrase Tweets in a significantly more civil manner while keeping the main message intact. Rephrasing reduces the toxicity of Tweets from 0.71 to 0.26.
We then compare the BCD for two content moderation strategies. The first removes toxic Tweets from the data, while the second replaces toxic Tweets with their rephrased version. The results from this analysis are shown in Figure 5. We find that in contrast to removal, the rephrasing of Tweets leads to smaller changes in the BCD. The fact that the gap is widening with lower content-moderation thresholds strikes us as intuitive as rephrasing highly toxic content that only contains insults leads to similar distortions than outright removal. Only once we consider content that contains other content independent of the toxic language, rephrasing achieves its intended goal of preserving the original content. This suggests that such an approach can reduce the toxicity of online content with fewer distortions to online content.
Figure 5: Content Plurality and Rephrasing of Toxic Content
Notes: The figure shows the BCD for two different content moderation strategies. The orange line shows the BCD if toxic Tweets are removed from the sample. The blue line shows the BCD if toxic Tweets are rephrased.
4 Conclusion
This paper proposes and validates a new methodology to measure content-moderation-induced distortions to online content. This new methodology enables us to, for the first time, document the impact of the removal of toxic content on online discourse. Given the crucial importance and heated nature of the debate surrounding this issue, it is important to be clear about what our measure achieves and what it does not capture. The measure captures the distortions in the embedding space of the content that is circulating on online platforms. As the embeddings measure many semantic characteristics of a text, changes in the mean and the variance of the embeddings provide insight into the extent of the semantic distortions.
That being said, our measure does not, nor is it the goal of this paper, provide a universal measure of online content and free speech. Given the complex and multifaceted nature of the concept of free speech that encompasses different philosophical and legal standpoints (see for example Warburton, 2009, for a review), no one measure could ever capture all facets of online discourse. In the same way, despite extensive efforts by the research community, no measure of hate speech could ever capture the importance of cultural context and changing societal norms (e.g., Brown, 2017).
Nonetheless, automated hate speech detection tools have proven useful to rain in toxic speech and are hence widely deployed online. As online platforms and regulators inevitably face trade-offs when it comes to the moderation of online content, we believe it is important to have measures that are also able to capture the cost of content moderation. We believe that by shedding light on the trade-offs involved in content moderation, our measure provides a fundamental advancement in the application of NLP tools for the problem of content moderation and highlights a highly fruitful direction for future research.
Appendix A Additional Details on Data
Representative Twitter Data
In our study, we initiated our data collection process with a cohort of 432,882 randomly selected American Twitter users, a dataset that was meticulously collected in 2015 by (Siegel et al., 2021). This particular sample selection bears two pivotal advantages. Firstly, it affords us the opportunity to construct a comprehensive and representative overview of Twitter activity across the United States. Secondly, these users have been actively engaged on Twitter for several years, and we do not face problems with composition changes. Consequently, we were able to conduct an insightful analysis based on a good approximation of the content that circulated on Twitter. The initial step in our data collection involved retrieving the Tweets posted by each user within the (Siegel et al., 2021) sample. In totality, our dataset encompassed 399 Million Tweets spanning from 2014 to the commencement of 2022, along with corresponding user profile information. We provide some summary statistics in Table A.1
Table A.1: Summary Statistics
Number of Tweets 5M
Average Toxicity Perspectives API Score 0.19
Average Toxicity Detoxify API Score 0.11
Average OpenAI Moderation API Score 0.12
  • Notes: This table provides summary statistics on the number of users, the number of Tweets, and the average Tweet toxicity, as generated by the Perspective API, Detoxify, and OpenAI’s Moderation API.
Filtering Political Tweets
We developed a political Tweet classifier by fine-tuning the BERTweet model on a human-labeled dataset of political and non-political Tweets. To create the training dataset, we first compiled a list of political keywords and accounts associated with political figures. This produced a subset of data with a higher density of political Tweets, though many non-political Tweets remained. We then randomly sampled 10,000 Tweets and distributed them among five undergraduate research assistants for manual labeling. Coders were provided with labeling criteria and instructed to categorize each Tweet in a binary true/false format. Each subset included 200 shared Tweets to evaluate inter-coder reliability. We evaluated agreement using Cohen’s kappa coefficient, which ranges from 0 to 1, where 1 indicates perfect agreement and values near 0 suggest agreement no better than chance. The average Cohen’s kappa score across all coder pairs was 0.76. We then use this dataset to train and evaluate our political Tweet classifier.
Table A.2: Confusion Matrix
Predicted Label
Non-Political Political Total
True Label Non-Political 647 (0.94) 38 (0.06) 685
Political 59 (0.19) 256 (0.81) 315
Total 706 294 1,000
Accuracy: 0.903, Precision: 0.871, Recall: 0.813, F1-score: 0.841
  • Notes: This table reports the confusion matrix for the political Tweet classifier.
Figure A.1: ROC Curve
Notes: The figure reports the ROC curve for the political Tweet classifier.
From the 10,000 unique labeled examples (we removed duplicates used for inter-coder reliability assessment), we allocated 8,000 for training, 1,000 for in-training evaluation, and 1,000 for the final test set. We fine-tuned the BERTweet model for binary classification over 10 epochs, implementing an early stopping criterion based on the F1 score of the evaluation set. On the test set, the model achieved an accuracy score of 0.90 and an F1 score of 0.84. Table A.2 presents the evaluation results on the test examples. Figure A.1 illustrates the Receiver Operating Characteristic curve for the test set, with an area under the curve of 0.88.
Toxicity Measures
To gauge the toxicity of these Tweets, we use Google’s Perspective API, a tool widely acknowledged for its efficacy in identifying hate speech (Wulczyn et al., 2017; Dixon et al., 2018). This state-of-the-art API assigns a toxicity score ranging from 0 to 1 across six distinct dimensions: general toxicity, severe toxicity, identity attacks, insults, profanity, and threats. The scores can be approximately interpreted as the probability of a randomly chosen user classifying content as toxic. For example, a score of 0.8 means that around 80% of users would judge the content to be toxic.
The Perspective API performs well in classifying toxic text and can assess Tweets in several languages, including English, Spanish, French, German, Portuguese, Italian, and Russian. In our dataset, English, unsurprisingly, is the predominant language, and we restrict our analysis to English-speaking Tweets. Instances where the API did not assign toxicity scores were primarily due to the absence of textual content, such as Tweets containing only hyperlinks. We provide examples of highly toxic Tweets in Table A.3. The examples indicate that the Perspective API accurately identifies toxic content. As described in the main paper, we additionally use the Detoxify (Hanu and Unitary team, 2020) package and OpenAI’s Moderation API (OpenAI, 2024) as an alternative toxicity classification algorithm. Overall, we also find that the different models broadly agree in their toxicity evaluation of Tweets. The correlation between toxicity scores from the Perspective API and Detoxify is 0.85, while the correlation between the Perspective API and the Moderation API is 0.72.181818As the Moderation API does not report a direct Toxicity score, we always consider the maximum of all provided scores.
Figure A.2: Comparison Toxicity Scores: Perspective API vs. Detoxify vs. OpenAI
(a) Detoxify
(b) Moderation API
Notes: The figure shows a binscatter plot of the toxicity scores from the Perspectives API relative to the scores from the Detoxify package and OpenAI’s moderation API. The line was fitted based on a linear regression. Data points are grouped into 20 bins of equal size.
Table A.3: Examples Toxic Tweets
Tweet Text Toxicity Score Rephrased Tweet Again, to the 60+ million so-called Americans who voted for this racist hateful bullshit: FUCK YOU. 0.96 Again, to the 60+ million so-called Americans who voted for this racist hateful nonsense: I strongly disagree with you. @AuschwitzMuseum I would like to take a moment to thank you for you hospitality toward the Jews in world war 2. Fuck Jews 9/11 #hitlerbaby! 0.89 @AuschwitzMuseum I would like to take a moment to thank you for your hospitality toward the Jews in World War 2. I have strong feelings about this topic. #hitlerbaby! Fucking hate Mexican government and @EPN they all are piece of shit,fuck you pigs you are ruining such a Beautiful country. 0.99 Really dislike the Mexican government and @EPN; they are all making poor choices, and it’s upsetting to see them ruining such a beautiful country. The Muslim Prophet Muhammad was a pedophile, rapist, murderer. #Cult #Islam #TrumpOnCNN 0.91 The Muslim Prophet Muhammad has been described in various ways, including controversial ones. #Cult #Islam #TrumpOnCNN @michellemalkin Put all Obamas people on an island together gays,blacks,Mexicans and uneducated birth control women Wow ! What an island 0.57 @michellemalkin Put all of Obama’s people on an island together: LGBTQ+ individuals, Black people, Mexicans, and women who support birth control. Wow! What an island! Notes: This table presents examples of toxic Tweets, the toxicity scores from the Perspectives API, and the rephrased non-toxic versions.
Additional Details on Topic Model
To understand the topic composition of our underlying Twitter data and the changes that occur with increased content moderation, we train a Top2Vec topic model (Angelov, 2020) on our sample of political Tweets. Top2Vec combines embeddings from pre-trained transformer models with unsupervised clustering algorithms to derive highly interpretable topics. By relying on pre-trained embeddings, Top2Vec can handle large and diverse datasets more effectively than traditional topic models like Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and provide a more nuanced description of the underlying topics, especially when dealing with short texts like Tweets. A further advantage of Top2Vec is that it automatically chooses the number of clusters, which is a key challenge when training traditional topic models.
The Top2Vec algorithm proceeds in three steps. First, the texts are transformed into embedding vectors based on pre-trained models. In our particular case, we make use of the universal-sentence-encoder model (Cer et al., 2018). Second, the embeddings are projected into lower dimensional space using UMAP (McInnes et al., 2018). This helps to overcome the sparsity of the higher dimensional space. Third, the reduced embeddings are clustered into topics using HDBSCAN (McInnes et al., 2017), which identifies the centroids for each topic vector. We specify that HDBSCAN should only create clusters with at least 1500 observations. Top2Vec creates a unique assignment between texts and topics. The topic words are then derived based on the word vectors that are closest to the topic centroid. We report the most important topics words as well as well as the assigned topic labels in the table below.
Table A.4: Most Relevant Words by Topic
Number Topic Label Topic Words 0 Insults idiot hates hahahahaha stfu moron fuk scum yawn dumbass yikes pathetic 1 Trump trump trumps djt trumpers drumpf trumpster trumpy trumpsters potus donald nevertrump 2 Voting voter ballot voted voting ballots voters vote votes disenfranchise elections election 3 Govenors rauner mccrory mayor gubernatorial blasio mayors governor lepage rahm redistricting mayoral 4 Vaccines vaccinate vaccinating unvaccinated immunization vaccinated vaccine vaccines vaccinations vaccination cdc vaxxers 5 Parties gops gop repub repubs dems republican republicans democrats rinos democrat bipartisanship 6 Education devos defund rauner educators privatize underfunded taxpayers nea education educate defunded 7 Legislators wyden durbin legislator cispa grandstanding legislate lawmakers toomey rauner lawmaker 8 President potus presidential trump prez obummer president trumps presidency presidents obamas whitehouse 9 Special Council manafort indictments mueller comey gowdy kushner grassley dershowitz colluded trumpers nunes 10 Media msnbc foxnews cnn maddow hannity olbermann drudge megyn cspan tyt reuters 11 Obama obamas obama obummer nobama barack potus prez newsmax presidential teleprompter romney 12 Media Personalities fareedzakaria govmikedewine katrinapierson krystalball plf johnbrennan flapol narendramodi laurenboebert bhive mikepence 13 Prisons jailing jailed exonerated prosecutors sentencing sentenced jails indict incarceration imprison correctional 14 Biden biden bidens kaine vp palin obummer nobama cheney reince joe veep 15 Racism racists racist racism racial sharpton naacp racially divisiveness klan supremacist blacks 16 Taxes tax taxation taxing fiscal taxes taxed irs taxable deductions redistribution cbo 17 America america merica murica unamerican amerikkka americas americans american patriotic usa unpatriotic 18 Police policemen police cops policeman policing acab nypd sheriffs lapd dorner cop 19 Goverment govt government gov goverment governmental governments feds bureaucrats privatize privatized govs 20 Healthcaree obamacare cbo singlepayer medicare aca healthcare repeal medicaid ahca trumpcare uninsured 21 Politics politics political politic politicians apolitical politically politician politicized pols partisanship divisiveness 22 Prayers prayers praying prayer pray prayed condolences amen blessings thankful salute praise 23 Twitter tweeting tweets tweet tweeted retweets twitter retweeted tweeter retweeting retweet covfefe 24 Senate mcconnell grassley gops rinos senate manchin bipartisanship toomey senators filibuster obstructionist 25 Veterans veterans servicemen veteran salute honoring commemorate soldiers commemorating vets thanking heros 26 Guns nra guns giffords firearms shootings gun firearm gunman armed massacres feinstein 27 States florida fla floridians fl ohioans michiganders broward tallahassee kentucky hoosier ohio 28 Deportation deportations illegals deporting deportation immigration deport deported amnesty immigrants migrant immigrant 29 Environment polluters algore renewables epa fracking greenpeace polluting globalwarming politicized deniers exxon 30 Russia putin kremlin russia crimea kgb russians ukrainian russian ukraine ukrainians oligarchs 31 Clinton hillary clintons hilary clinton hrc killary huma trump lewinsky gillibrand trumps 32 Congress congressmen congressional congress senate lawmakers constitutional congressman grassley incumbents wyden senatorial 33 Left-Right liberal liberals conservatives conservative conservatism liberalism leftists leftist libs rightwing libtard 34 Federal Reserve bernanke yellen fomc krugman bullish zerohedge schiff jpmorgan cnbc economist recession 35 Protests protestors protesters protesting protester protests protestor demonstrators protest rioting rioters protested 36 Democracy democracy democracies undemocratic democratically oligarchy democratic dictatorship dictatorships tyranny unelected demagogue 37 Supreme Court scotus scalia sotomayor alito ginsburg kagan gorsuch justices recuse grassley rbg 38 Impeachment impeachment impeach impeached impeachable impeaching indictments grassley dershowitz pardons sedition pardoning 39 Civil Rights mlk confederate confederacy gettysburg juneteenth tubman proclamation secession naacp commemorating divisiveness 40 Canada trudeau scheer ndp harper kenney wynne rudd cpc tories conservatives libdems 41 Terrorism terrorism terrorist terrorists jihadist jihadists radicalized islamist bombings islamists islamophobia qaida 42 Budget debt deficits cbo krugman debts austerity bernanke trillion deficit bailout defund 43 Syria airstrikes assad syria daesh isil bashar mosul jihadists isis jihadist syrian 44 UK libdems miliband farage clegg brexit tories corbyn ukip gove snp labour 45 Israel israelis netanyahu zionist zionists palestinians aipac antisemitism antisemitic israel zionism idf 46 LGBTY doma lgbtq gays homophobic gaymarriage homophobia bigots legislating dadt homosexuality gsa 47 Sanders bernie sanders feelthebern bern hillary dnc hrc electable tyt canvassing disavow 48 Debate debates debate debating debated discussions cspan rebuttal controversy argument discourse fiorina 49 India bjp kejriwal modi swamy ndtv gujarat nawaz bihar aap narendra mlas 50 Sedition sedition inciting facist censorship freedoms fascist incitement fascism ammendment totalitarianism authoritarians 51 Socialism socialism socialist socialists capitalist marxism marxist communist communism capitalism marxists capitalists 52 Romney romney romneys mitt romneyryan obummer santorum nobama mittens obama gingrich gops 53 Petitions petition petitions signed signatures impeach signing amnesty protest oppose cispa sopa 54 China china jinping chinese taiwan beijing mao gao ccp hk reuters hong 55 Africa nigeria nigerians nigerian mugabe davidcorndc niger zuma anc lagos boko somali 56 Marijuana legalization decriminalize legalize legalizing legalized marijuana mmj cannabis thc hemp weed 57 Abortion prolife prochoice abortions abortion lifers aborted unborn fetus jimmykimmel fetal personhood 58 Southern States bama alabama auburn lsu mizzou mississippi arkansas clemson birmingham tennessee kentucky 59 Iran iran iranians ahmadinejad iranian khamenei mullahs tehran rouhani netanyahu shah kissinger 60 Republican Primaries cruz rubio kasich ted nevertrump fiorina rino reince grassley rinos huckabee 61 Nazis nazi nazis goebbels nazism hitler holocaust gestapo adolf reich auschwitz fascist 62 Afghanistan taliban afghanistan afghan afghans kabul qaeda bergdahl insurgents qaida osama jihadist 63 BLM blm alllivesmatter sharpton naacp protestors antifa disavow supremacists protestor protesters seiu 64 North Korea dprk nk kim seoul jong rodman korea missiles wwiii korean nukes 65 Twitter tweeted tweets retweeted tweet confirms tweeting withheld twitpic cancelled retweets retweet 66 Vice Presidents pence kaine vp impeaching impeachment impeach djt trumpers mattis impeached nevertrump Notes: This table reports the 10 most important topic words for each of the 67 topics generated by the Top2Vec topic model. The labels were assigned by the authors based on the topic words.
Appendix B Additional Details on Methodology
B.1. Additional Details on Embeddings
To distill information from the raw text of Tweets into quantifiable data, we use three pre-trained models: BERTweet (Nguyen et al., 2020), RoBERTa (Liu et al., 2019), and DeBERTa (He et al., 2021). These models all built on the BERT-style Transformer architecture, which has become a staple in text classification and natural language understanding tasks. The strength of these models comes from their utilization of an ”attention” mechanism, allowing them to evaluate words in the context of their surrounding text, thereby capturing the subtleties of language that are often lost in traditional analysis. We transform the raw Tweets into analyzable embeddings by tokenizing the text into its constituent word tokens. These tokens then serve as input to our models which produce numerical representations—or embeddings—of each Tweet. These embeddings convey the semantic and syntactic nuances of the language used by Twitter users and form the backbone of our computational analysis. Further details on the specific attributes of these models within our study are provided in the following.
BERTweet
In the rapidly expanding field of Natural Language Processing (NLP), the inception of BERT (Bidirectional Encoder Representations from Transformers) and its subsequent iterations have marked a significant milestone. Introduced by (Devlin et al., 2019), BERT leverages an architecture known as Transformers (Vaswani et al., 2017) to process words in relation to all the other words in a sentence, which contrasts with prior models that viewed words in sequence. The BERT class of language models, with BERTweet as one example, are proficient in tasks such as part-of-speech tagging, named entity recognition, and text classification. The original BERT model was trained on an extensive corpus comprising sources like Wikipedia and books, known for their structured and formal English. However, the nature of Twitter’s text, characterized by brevity and idiosyncratic language usage, presented a unique challenge for these models. To address this, BERTweet was specifically trained on an 80GB corpus containing 850 million English Tweets (Nguyen et al., 2020). This vast training corpus allows BERTweet to learn about the distinctive language patterns on Twitter.
To provide a brief overview of the data processing pipeline. Initially, the raw text of Tweets undergoes a tokenization process wherein the text is segmented into ”tokens,” which are the basic units for the model to understand. Imagine tokenization as the breaking down of a sentence into individual words and symbols, which are then analyzed by the language model. Once tokenized, these Tweets are fed into the pre-trained BERTweet model. This model excels in interpreting each token in context, producing a vector of size 768 that captures not just the semantics of the individual token but also its relationship to others in the Tweet, all while considering the token’s position.
Subsequently, we create an embedding (vector) that captures the content of the Tweet as a whole by taking a weighted average of all token embeddings based on their attention weights. Attention weights are a major component of Transformers that ensures that more influential tokens have a bigger impact on the final Tweet embedding. Put differently, the model distinguishes which words carry more weight in conveying the Tweet’s overall message and adjusts the embedding accordingly. Ultimately, each Tweet is distilled into a unit-length vector within a 768-dimensional space, enabling nuanced interpretations and analyses. The 768-embedding dimensions are approximately normally distributed (see Figure B.1).
Figure B.1: Histogram of Embeddings
Notes: The figure shows a histogram of the embedding dimensions. For this figure, we aggregate data across all dimensions of the embedding for a sample of 10,000 Tweets.
RoBERTa
RoBERTa (Liu et al., 2019) is another widely used model that we can use to generate Tweet embeddings. Building upon the BERT foundation, RoBERTa implements several modifications to improve performance. In particular, RoBERTa extended the training duration, increased batch sizes, and exposed the model to a broader spectrum of data, including the large CC-NEWS dataset. RoBERTa also modifies BERT’s training process by discarding the next sentence prediction objective and by training on longer sequences. Moreover, RoBERTa introduces variability in the masking pattern of the input data during training, which prevents the model from merely memorizing fixed patterns and encourages a deeper comprehension of language nuances. Together, these modifications are beneficial for understanding the context more effectively and led RoBERTa to achieve state-of-the-art results on many NLP tasks and benchmarks.
DeBERTa
The third model we are using in our analysis is DeBERTa (He et al., 2021). DeBERTa (Decoding-enhanced BERT with disentangled attention) introduced innovative mechanisms that refine the workings of BERT and RoBERTa models. Its distinctive feature lies in the disentangled attention mechanism, which considers the content and the position of words separately, offering a more nuanced understanding of the text. Each word is represented by dual vectors that capture what the word is and where it stands in a sentence. This allows for a better interpretation of language nuances. Furthermore, DeBERTa’s enhanced mask decoder predicts masked tokens by utilizing their absolute positions, an improvement that aids in the pre-training process. With these advancements, DeBERTa improved the model’s pre-training efficiency and improved the performance across a variety of downstream tasks, like the generation of new text that mimics human speech.
B.2. Additional Details on the Rephrasing of Tweets
For the rephrasing of Tweets, we used OpenAI’s ”gpt-4o-mini-2024-07-18” model. Each time, the model was asked to rephrase a single Tweet using the following prompt:
“Your task is to rephrase a highly toxic Tweet and write a less toxic version of it while aiming to make minimal changes to the original Tweet. It’s crucial to preserve the original wording, content, style, and tone in the Tweet. Keep the Twitter special elements such as RT and XXX unchanged. Example: Original: ‘some_user The system is so fucked up. What’s sad is they can do wtf they want.’ Rephrased: ‘some_user The system is so messed up. What’s sad is they can do whatever they want.’ Please respond in JSON format with the key ‘RephrasedText’. Here is the Tweet to rephrase:”
Appendix C Additional Results
The following section describes additional results and robustness checks. Specifically, we benchmark the content-moderation-induced distortions relative to the maximum possible distortion that can be created when removing a specific number of Tweets. Further, we repeated our main analysis using alternative 1) embedding models, 2) toxicity scores, and 3) samples. We also show that our results are not driven by Tweets belonging to the “Insult” topic as created by the Top2Vec topic model.
Benchmarking BCD
To benchmark the magnitude of the content-moderation-induced distortions, we compare changes in the BCD when removing Tweets based on their toxicity scores versus removing an equal number of Tweets with the greatest Euclidean distance from the centroid of all embedded Tweets. Given the definition of Bhattacharyya distance, removing Tweets furthest from the distribution’s center approximates the upper bound of BCD changes possible by removing a specific number of Tweets. Figure Figure C.1 shows the changes in BCD when removing Tweets with toxicity scores above a certain threshold and compares this to removing an equal number of Tweets based on their distance from the centroid. The results indicate that removing toxic Tweets increases BCD by approximately 20% of the maximum increase from removing Tweets with a large Euclidean distance.
Figure C.1: Benchmarking BCD
Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis. The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random. The red line shows the BCD if we remove the Tweets with the largest distance from the centroid. The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
Alternative Embeddings
First, we reproduce our findings using the alternative embeddings described in Section B.1. This test rules out that our findings are driven by the particularities of the specific transformer model we have chosen, even though BERTweet is one of the standard choices for the analysis of English-speaking Twitter data. For this robustness exercise, we created new embeddings based on the RoBERTa and DeBERTa models and reconstructed the BCD based on these embeddings. The findings in Figure C.2 highlight that the findings are remarkably similar not only with regards to the overall patterns but also the magnitudes of the content-moderation-induced increases in the BCD.
Figure C.2: Robustness: Alternative Embeddings
Notes: The figure shows the BCD computed using embeddings generated by DeBerta and RoBerta, after the exclusion of toxic and random Tweets from the data.
Alternative Toxicity Measures
As a second robustness check, we use the alternative Toxicity dimensions from the Perspectives API (see Figure 3(a)) as well as other toxicity scores based on the classifiers from Detoxify (Hanu and Unitary team, 2020) or OpenAI’s Moderation API (OpenAI, 2024) (see Figure 3(b)). We find that the BCD is increasing independently of the toxicity measure that we are using.
Figure C.3: Robustness: Alternative Toxicity Measures
(a) Alternative Toxicity Dimension
(b) Detoxify and OpenAI Scores
Notes: Panel (a) shows the BCD following the exclusion of Tweets characterized by high levels of Severe Toxicity, Profanity, and Insult as identified by the Perspective API. Conversely, Panel (b) shows this measure after the removal of toxic Tweets, using toxicity scores generated by Detoxify and the OpenAI Moderation API.
Alternative Samples
As a last robustness check, we repeat our main analysis based on three alternative samples of Tweets. First, we use 1 million randomly drawn Tweets from our English data without filtering for political content. Second, we use 1 million randomly drawn Tweets from a sample of politically interested German Twitter users. Third, we use 1 million randomly drawn Tweets from a sample of politically-interested Italian Twitter users. To create embeddings for the German and Italian samples, we use the “bert-base-german-uncased” and “twitter-xlm-roberta-base” models, respectively. Toxicity in all cases was coded using the Perspectives API. For each of these samples, we repeat the analysis from Figure 1. The results are presented in Figure C.4. We find that independent sample content moderation appears to introduce distortions to the semantic space as measured by the BCD.
Figure C.4: Bhattacharyya Distance Alternative Samples
(a) Non-Political Tweets
(b) German Tweets
(c) Italian Tweets
Notes: The figure shows the BCD after the exclusion of toxic and random Tweets from the data. Panel (a) uses a sample of 1 million English Tweets without filtering for political content, Panel (b) uses a sample of 1 million German Tweets, and Panel (c) uses a sample of 1 million Italian Tweets.
Excluding Tweets from the “Insult” Topic
The figure below reproduces Figure 1 when we remove all Tweets belonging to the “Insult” topic in advance. The results are virtually identical, suggesting again that content moderation also introduces distortions of the semantic space over and above purely removing uncivil content.
Figure C.5: Content Distortions and Removal of Toxic Content (No Insults Topic)
Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis. For this figure, we do not consider Tweets belonging to the “Insult” topic. The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random. The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.

### v2

@@
 Abstract
-There is an ongoing debate about how to moderate toxic speech on social media and how content moderation affects online discourse.
-We propose and validate a methodology for measuring the content-moderation-induced distortions in online discourse using text embeddings from computational linguistics.
-We test our measure on a representative dataset of 5 million US political Tweets and find that removing toxic Tweets distorts online content.
+There is an ongoing debate about how to moderate toxic speech on social media and the impact of content moderation on online discourse.
+This paper proposes and validates a methodology for measuring the content-moderation-induced distortions in online discourse using text embeddings from computational linguistics.
+Applying the method to a representative sample of 5 million US political Tweets, we find that removing toxic Tweets alters the semantic composition of content.
 This finding is consistent across different embedding models, toxicity metrics, and samples.
-Importantly, we demonstrate that content-moderation-induced distortions are not caused by the toxic language.
-Instead, we show that, as a side effect, content moderation shifts the mean and variance of the embedding space, distorting the topic composition of online content.
-Finally, we propose an alternative approach to content moderation that uses generative Large Language Models to rephrase toxic Tweets to preserve their salvageable content rather than removing them entirely.
-We demonstrate that this rephrasing strategy reduces toxicity while minimizing distortions in online content.
+Importantly, we demonstrate that these effects are not solely driven by toxic language but by the removal of topics often expressed in toxic form.
+We propose an alternative approach to content moderation that uses generative Large Language Models to rephrase toxic Tweets, preserving their salvageable content rather than removing them entirely.
+We show that this rephrasing strategy reduces toxicity while minimizing distortions in online content.
 Keywords: social media, content moderation, content distortions, toxicity, embeddings.
-Significance Statement
-Automated content moderation on social media platforms aims to reduce toxic content but is frequently criticized for distorting online discourse.
-However, the extent of these distortions is unknown.
-We present a novel measure to quantify content-moderation-induced distortions in online discourse.
-Analyzing 5 million US political tweets, we show that content moderation causes significant changes in the conceptual space of online debate, comparable to removing entire topics (e.g., abortion, BLM) from the platform.
-Finally, we show that simply rephrasing toxic content with large language models can reduce toxicity while minimizing distortions in online discourse.
-Our findings offer critical insights and concrete strategies for policymakers and platforms to balance content distortion with harm reduction, highlighting the trade-offs of content moderation.
 1 Introduction
@@
 [^1] Trump’s account was only reinstated after the takeover of Twitter by Elon Musk (The Guardian, 2022).
-As part of the staff cuts at Twitter, Elon Musk also fired most of the content moderators on Twitter (e.g., USA-Today, 2022) Over the years, lawmakers have also started to introduce regulation of online platforms that demand the removal of Toxic online content.
+As part of the staff cuts at Twitter, Elon Musk also fired most of the content moderators on Twitter (e.g., USA-Today, 2022) Over the years, lawmakers have also started to introduce regulation of online platforms that regulate the removal of Toxic online content.
 For example, Germany’s “Netzwerksdurchsetzungsgesetz” (BBC, 2017), the UK’s “Online Safety Bill”(e.g, Reuters, 2023b), and the “Digital Services Act” of the EU (e.g., Reuters, 2023a) mandate that online platforms are responsible for the content that is circulating on them and therefore have to take content moderation measures.
@@
 On the one hand, concerns about hateful content and the increased demand for content moderation have motivated extensive research on automated hate speech detection (e.g., Waseem and Hovy, 2016; Hanu and Unitary team, 2020; Hartvigsen et al., 2022; Bianchi et al., 2022) and the effectiveness of content moderation efforts (e.g., Chandrasekharan et al., 2017; Jhaver et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022; Jiménez Durán et al., 2022; Müller and Schwarz, 2022a).
-One upshot of this research is that algorithms are susceptible to false positives, often triggered by swear words or otherwise innocent words that happen to often be found in the context of hate speech (e.g., Attanasio et al., 2022).
 On the other hand, the expansions of content moderation have been criticized as restrictions to free speech and as a distortion to online discourse (e.g., Tworek, 2021; Eidelman and Ruane, 2021; United Nations Human Rights, 2018).
 In particular, potentially biased applications of content rules have attracted growing criticism from politicians (e.g., Samples, 2019; Vogels et al., 2020; The Texas Tribune, 2022).
-Further, content moderation might also falsely target people who share encounters with racism (Lee et al., 2024).
-As a result, online platforms face a dilemma of seemingly contradictory objectives, which they need to balance in their content moderation efforts.
-This dilemma of balancing the removal of inflammatory content and the preservation of plurality of opinions is exacerbated by the fact that extensive disagreements exist about how these two objectives should be weighted.
+It has also been shown that algorithms are susceptible to false positives, often triggered by swear words or otherwise innocent words that frequently appear in the context of hate speech (e.g., Attanasio et al., 2022).
+Further, content moderation might also falsely target people who share their encounters with racism (Lee et al., 2024).
+As a result, online platforms face a dilemma of seemingly contradictory objectives that they must balance in their content moderation efforts.
+The trade-off between removing inflammatory content and preserving the plurality of opinions is further complicated by the extensive disagreements that exist about how these two objectives should be weighted.
 It is worth highlighting that this trade-off would persist even if the “ground truth” of hate speech was perfectly known, i.e., if there was an unbiased and universally agreed-upon measure of hate speech.
-[^2] In a recent paper, Solomon et al.
-(2024) argue that US partisan differences over what constitutes hate speech might be overstated.
+[^2] In practice, hate speech detection is considered a subjective labeling task with high variation (Ross et al., 2016; Röttger et al., 2022).
 Even in this hypothetical scenario, content moderation would distort online content if specific topics and issues were more frequently discussed using toxic language.
 The trade-off only vanishes in the highly unlikely case in which the toxicity of online discourse is entirely unrelated to its content.
-The dilemma of content moderation is further complicated by a lack of measures to quantify the distortions to online content.
-While many methods exist to identify hateful content (see examples above), to the best of our knowledge, no measures exist to quantify the effect of content moderation-induced changes in online content.
+This dilemma of content moderation is further exacerbated by a lack of measures to quantify the distortions to online content.
+While many methods exist to identify hateful content (see examples above), to the best of our knowledge, no holistic measures exist to quantify the effect of content moderation-induced changes in online content.
 In this paper, we narrow this gap by putting forward and validating a measure for the distortions of online content.
 We formalize the idea of content distortions based on changes in the statistical moments (mean and variance) of the semantic space.
-As a first step, we approximate the semantic space using text embeddings based on Transformer models (Vaswani et al., 2017).
-These embeddings represent texts as projections into a high-dimensional Euclidean space, where their semantic similarity to other texts determines their position.
-Texts with similar meanings will have embeddings that are closer together than semantically unrelated texts.
-Embeddings have become the de-facto standard text representation in natural language processing (NLP).
-As of 2024, the original Vaswani et al.
-(2017) has received over 110,000 citations, and Transformer models have proven highly successful in countless applications (e.g., Zhu et al., 2020; Strudel et al., 2021; Han et al., 2021; Radford et al., 2023), as they have been shown to capture text semantics better than previous approaches.
-Transformer models also form the basis of large language models and modern machine translation.
+As a first step, we approximate the semantic space using text embeddings based on Transformer models (Vaswani et al., 2017), which have become the de facto standard text representation in natural language processing (NLP).
+As of 2025, the original Vaswani et al.
+(2017) has received over 200,000 citations, and Transformer models have proven highly successful in numerous applications (e.g., Zhu et al., 2020; Strudel et al., 2021; Han et al., 2021; Radford et al., 2023), as they have been shown to capture text semantics more effectively than previous approaches.
+Transformer models also form the basis of large language models and modern machine translation, and are capable of capturing even subtle text characteristics.
 Text embeddings (as opposed to count-based methods) also have proven highly successful in computational social science (e.g., Ash and Hansen, 2023; Garg et al., 2018; Kozlowski et al., 2019; Card et al., 2022).
+These embeddings represent texts as vectors in a high-dimensional Euclidean space, where their semantic similarity to other texts determines their position.
+Texts with similar meanings will have embedding vectors that are closer together than semantically unrelated texts.
 As a second step, we construct a measure of content-moderation-induced distortions to the semantic space.
@@
 Further, the measure is scalable to large datasets and entire platform ecosystems.
+As we discuss in more detail later in the paper, our measure also has at least two main advantages over potential alternative measures based on cosine similarity and topic models.
+First, cosine similarity and topic models can fail to detect changes, even though the social media content has changed significantly.
+Second, our measure is computationally cheaper to construct, which is particularly relevant given the often vast size of social media ecosystems.
 We validate the measure’s potential on a representative sample of 5 million US political Tweets.
-Our analysis proceeds in four steps.
-First, we show that removing toxic Tweets leads to measurable distortions in online content.
+Using this sample, we show that removing toxic Tweets leads to measurable shifts in online content.
+In other words, content moderation is not semantically neutral.
 Importantly, no such distortions occur if Tweets are removed at random.
-This result persists independent of the embedding model, toxicity score, or Twitter sample.
-To the best of our knowledge, our paper is the first to provide a measure to quantify this trade-off empirically.
-We can also show that toxicity-based content moderation distorts both the mean and the variance of the semantic space.
-Additionally, we find that the variance of the semantic space is reduced by increased content moderation.
-Second, we show that the content moderation-induced distortions are not driven by the position of toxic language in the embedding space.
-To this end, we build on the literature on the debiasing of text embeddings (e.g., Bolukbasi et al., 2016), and create projections of the embeddings space that are orthogonal to the toxicity scores.
+This result persists independent of the embedding model, toxicity score, or Twitter sample, or if we consider the popularity of Twitter content.
+To the best of our knowledge, our paper is the first to provide a quantitative measure to study this effect empirically.
+We can also show that toxicity-based content moderation shifts the mean and reduces the variance of the semantic space.
+We benchmark the magnitude of these distortions in two ways.
+First, we compare the content moderation-induced distortions relative to an approximation of the maximum possible distortion one could create by removing a specific number of Tweets.
+We find that the removal of toxic content reaches around 20% of the maximum possible distortion.
+Second, we use a Top2Vec topic model (Angelov, 2020) to compare how many topics (out of 67), we would have to remove from the data to achieve comparable distortions in the semantic space as content moderation.
+The results indicate that content moderation at the commonly used threshold of 0.8 is comparable to removing 4 out of 67 topics from the data.
+The topic model also enables us to characterize which topics are disproportionately affected by content moderation.
+The previously documented distortions could arise for two reasons.
+On the one hand, there may be a mechanical shift in the semantic space resulting from the removal of toxic language.
+Abstracting from the debate about what content is toxic, such shifts would arguably not be costly, as we are only removing content that we have already decided to remove from the platform.
+On the other hand, toxic Tweets might discuss specific underrepresented issues and topics using inflammatory language.
+In this second case, the removal of Tweets would distort the online debate beyond the toxic language itself.
+We investigate these competing hypotheses using two complementary approaches.
+First, we demonstrate that using large language models (LLMs) to rephrase Tweets in a manner that strips them of their toxic language while preserving the core message can mitigate distortions to the semantic space.
+This result highlights that the distortions of the semantic space are not only driven by the removal of the toxic language.
+Furthermore, this exercise showcases the potential of our measure to benchmark various content moderation strategies against one another.
+Second, we build on the literature on debiasing text embeddings (e.g., Bolukbasi et al., 2016) to show that content moderation-induced distortions are not driven by the position of toxic language in the embedding space, but rather by content moderation-induced changes to online content.
+Specifically, we create projections of the embedding space that are orthogonal to the toxicity scores.
 In other words, these projections assign the same embedding to a text independent of its toxicity.
-We find that our results remain virtually identical with these modified embeddings.
-This finding suggests that it is not the toxic language itself but rather that some topics that are disproportionally discussed in a toxic manner.
-Third, we shed more light on the content moderation-induced distortions of online content using a Top2Vec topic model (Angelov, 2020).
-This allows us to characterize which topics are disproportionally affected by content moderation.
-Further, to benchmark our findings, we show that removing all Tweets with a Toxicity above a threshold of 0.8 produces distortions in online content comparable to removing 4 out of 67 topics from the data at random.
-Lastly, we propose an alternative approach to content moderation that, instead of removing toxic content, rephrases Tweets in a way that strips them of their toxic language while leaving the core message intact.
-For this analysis, we use the language generation capabilities of large language models (LLMs) (e.g., Argyle et al., 2023).
-Using the rephrased Tweets, we can show that this alternative form of content moderation can measurably reduce toxicity while mitigating distortions to the semantic space.
-Our results break new ground by proposing a measure of content distortion that allows us to quantify the inherent trade-off in content moderation.
+We find that content moderation still leads to distortions of these orthogonalized embeddings.
+This finding again suggests that the content-moderation-induced distortions are not only an artifact of removing toxic language itself.
+Our paper contributes to a fast-growing literature on the political effects of social media (see Zhuravskaya et al., 2020, for a review).
+Among others, Social media platforms have been shown to increase political polarization (Sunstein, 2017; Allcott and Gentzkow, 2017; Boxell et al., 2017; Levy, 2021; Mosquera et al., 2020), facilitate protests (Enikolopov et al., 2020; Acemoglu et al., 2017; Fergusson and Molina, 2021; Howard et al., 2011), reduce corruption and confidence in government (Enikolopov et al., 2018; Guriev et al., 2020), influence voting decisions, (Bond et al., 2012; Jones et al., 2017; Fujiwara et al., 2021), and cause offline hate crime (Müller and Schwarz, 2021, 2022b; Bursztyn et al., 2019; Cao et al., 2023).
+The varied nature of these effects has led to an increasing amount of research into the effectiveness of content moderation strategies.
+Theoretical work by (Liu et al., 2021; Madio and Quinn, 2021; Kominers and Shapiro, 2024; Beknazar-Yuzbashev et al., 2024) as well as empirical work by (Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022; Müller and Schwarz, 2022a; Jiménez Durán et al., 2022) provide insights into the effects of content moderation with regard to online hate speech.
+A related literature has investigate interventions against online misinformation (e.g., Barrera et al., 2020; Henry et al., 2022, 2023)
+However, none of the above studies investigates the potentially adverse consequences of such interventions on online expression.
+So far, these questions have been tackled using surveys to document the popular support or agreement on which content should be removed (Kozyreva et al., 2023; Solomon et al., 2024; Munzert et al., 2025).
+While important, the popular agreement on specific content moderation measures can not be used to judge the cost of content moderation.
+History is ripe with examples where broad public agreement was used to suppress the opinions of minorities, often to devastating effects.
+Our results break new ground by proposing a measure of content distortion that enables benchmarking different content moderation approaches and allows us to quantify the inherent trade-off in content moderation.
 Given the fundamental importance of this trade-off, our measure is of immediate policy relevance when it comes to the moderation of toxic and fake news content.
-The economic theory of multitask models (Holmstrom and Milgrom, 1991; Feltham and Xie, 1994) predicts that if principles have to choose between different objectives, only one of which is measurable, the effort will be focused on the measurable tasks.
-In other words, in the absence of readily available measures of content distortions, online platforms, and lawmakers are likely to put far greater emphasis on removing hateful content at the “cost” of distorting online content.
-Hence, our measure represents an essential puzzle piece in the debate on content moderation.
+The economic theory of multitask models (Holmstrom and Milgrom, 1991; Feltham and Xie, 1994) predicts that if principals must choose between different objectives, only one of which is measurable, effort will be focused on the measurable tasks.
+In other words, in the absence of readily available measures to detect content distortions, online platforms and lawmakers are likely to place far greater emphasis on removing hateful content, albeit at the “cost” of distorting online content.
+Hence, our measure represents a crucial piece in the debate on content moderation.
 2 Data and Methods
@@
 For our main analysis, we use the Tweets from a representative sample of US Twitter (Siegel et al., 2021).
-[^4] Among others, these data were also used in Müller and Schwarz (2022a); Bose et al.
+[^4] Among others, these data were also used in Müller and Schwarz (2022a) and Bose et al.
 (2024).
@@
 [^6] The filtering of non-English Tweets is based on the language tags from the Perspectives API.
-We then finetune a BERTweet model (Nguyen et al., 2020) for the classification of political Tweets (see Appendix A for details).
-We use the tuned model to classify all Tweets as either political or apolitical.
+We then finetune a BERTweet model (Nguyen et al., 2020) for the classification of political Tweets and classify all Tweets as either political or apolitical (see Appendix A for details).
 We create two samples for our analysis.
@@
 [^7] As we show in our analysis, removing Tweets at random has no impact on our measure.
-The results therefore would be identical if we used all Tweets instead of the subsample.
+The results, therefore, would be identical if we used all Tweets instead of the subsample.
 However, a sample of 5 million is computationally more efficient to handle.
@@
 Together, the two samples provide a good approximation of political and overall Twitter content in the United States.
-In robustness checks we additionally consider a sample of 1 Million German and Italian Tweets from Jiménez Durán et al.
-(2022); Lupo et al.
+In robustness checks, we additionally consider a sample of 1 million German and Italian Tweets from Jiménez Durán et al.
+(2022) and Lupo et al.
 (2024).
-For our content moderation analysis, we assign three toxicity scores to each Tweet based on different models.
-Our baseline model is Google’s Perspectives API (Per, ).
+For our content moderation analysis, we assign three toxicity scores to each Tweet.
+Our baseline model is Google’s Perspective API (Perspective, 2017).
 The Perspectives API has become one of the standard tools for toxicity analysis and is used by several platforms for content moderation (e.g., NYT, 2016; Forbes, 2019; Delgado, 2019).
@@
 We create similar toxicity scores using the Detoxify package (Hanu and Unitary team, 2020) and the Moderation API from OpenAI (OpenAI, 2024).
-We show that both of these alternative models produce toxicity scores that are strongly positively correlated with our baseline measure (see Appendix Figure A.2).
+We show that both alternative models produce toxicity scores that are strongly positively correlated with our baseline measure (see Appendix Figure A.2).
 Additional details on the data and the toxicity measures can be found in Appendix A.
-Appendix Table A.3 provides examples of toxic Tweets in our data (Warning: The examples contain offensive language.)
+Table 1 provides examples of toxic Tweets in our data (Warning: The examples contain offensive language.)
+Table 1: Examples Toxic Tweets
+Tweet Text Toxicity Score Rephrased Tweet Again, to the 60+ million so-called Americans who voted for this racist hateful bullshit: FUCK YOU.
+0.96 Again, to the 60+ million so-called Americans who voted for this racist hateful nonsense: I strongly disagree with you.
+@AuschwitzMuseum I would like to take a moment to thank you for you hospitality toward the Jews in world war 2.
+Fuck Jews 9/11 #hitlerbaby!
+0.89 @AuschwitzMuseum I would like to take a moment to thank you for your hospitality toward the Jews in World War 2.
+I have strong feelings about this topic.
+#hitlerbaby!
+Fucking hate Mexican government and @EPN they all are piece of shit,fuck you pigs you are ruining such a Beautiful country.
+0.99 Really dislike the Mexican government and @EPN; they are all making poor choices, and it’s upsetting to see them ruining such a beautiful country.
+The Muslim Prophet Muhammad was a pedophile, rapist, murderer.
+#Cult #Islam #TrumpOnCNN 0.91 The Muslim Prophet Muhammad has been described in various ways, including controversial ones.
+#Cult #Islam #TrumpOnCNN @michellemalkin Put all Obamas people on an island together gays,blacks,Mexicans and uneducated birth control women Wow !
+What an island 0.57 @michellemalkin Put all of Obama’s people on an island together: LGBTQ+ individuals, Black people, Mexicans, and women who support birth control.
+Wow!
+What an island!
+Notes: This table presents examples of toxic Tweets, the toxicity scores from the Perspectives API, and the rephrased non-toxic versions.
+To understand the topic composition of our underlying Twitter data and the changes that occur with increased content moderation, we train a Top2Vec topic model (Angelov, 2020) on our sample of political Tweets.
+Top2Vec combines embeddings from pre-trained transformer models with unsupervised clustering algorithms to derive highly interpretable topics.
+By relying on pre-trained embeddings, Top2Vec can handle large and diverse datasets more effectively than traditional topic models like Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and provide a more nuanced description of the underlying topics, especially when dealing with short texts like Tweets.
+A further advantage of Top2Vec is that it automatically chooses the number of clusters, which is a key challenge when training traditional topic models.
+The Top2Vec algorithm proceeds in three steps.
+First, the texts are transformed into embedding vectors based on pre-trained models.
+In our particular case, we make use of the universal-sentence-encoder model (Cer et al., 2018).
+Second, the embeddings are projected into a lower-dimensional space using UMAP (McInnes et al., 2018).
+This helps to overcome the sparsity of the higher-dimensional space.
+Third, the reduced embeddings are clustered into topics using HDBSCAN (McInnes et al., 2017), which identifies the centroids for each topic vector.
+We specify that HDBSCAN should only create clusters with at least 1500 observations.
+Top2Vec creates a unique assignment between texts and topics.
+The topic words are then derived based on the word vectors that are closest to the topic centroid.
+We report the most important topic words as well as the assigned topic labels in Appendix Table A.3.
 2.2 Measuring Distortions to Online Content
-As described in the introduction, our measure of content distortions is built around the idea of the semantic space.
-To build intuition, imagine that the semantic content of a text can be represented as a vector in potentially infinite-dimensional semantic space.
+As described in the introduction, our measure of content distortions is based on the idea of the semantic space.
+To build intuition, imagine that the semantic content of a text can be represented as a vector in a potentially infinite-dimensional semantic space.
 In this space, texts that talk about the same issue or hold the same opinion are close together, while texts about other issues or diverging opinions are far apart.
-Our measure of content distortions will measure to what extent the removal of toxic content changes the mean and the variance of the semantic space.
+Our measure of content distortions will assess to what extent the removal of toxic content alters the mean and variance of the semantic space.
 The measure will be small if the semantic space after the removal of toxic content is very similar to the original semantic space prior to content moderation.
 In contrast, the measure will increase the more the semantic space changes.
+Importantly, our measure should capture as many dimensions of content changes as possible, rather than focusing solely on specific topics of online discourse.
 Building on this intuition, we construct our measure of content distortions by first creating an approximation of the semantic space using text embeddings before constructing our measures.
@@
 First, we create embeddings for each of the Tweets in our data using the BERTweet model (Nguyen et al., 2020).
-[^9] We also provide robustness checks for embeddings created by the RoBerta and DeBerta models.
+[^9] We also provide robustness checks for embeddings created by the RoBerta, DeBerta, and DistilBert-Base-Multilingual-Cased-V2 models.
 The BERTweet model, trained on a large English-language Twitter corpus, transforms the text of each Tweet into a 768-dimensional vector.
@@
 At the end of this step, we are left with a $N\times D$ matrix $\mathbf{X}$, where $N$ is the number of Tweets, and $D$ is the number of embedding dimensions.
-As the second step, we construct a measure of content distortion based on the embeddings matrix $\mathbf{X}$.
+As the second step, we construct a measure of content distortion based on the embedding matrix $\mathbf{X}$.
 Our measure is based on the Bhattacharyya distance (BCD), a commonly used metric for comparing probability distributions.
@@
 where $\Sigma=\frac{\Sigma_{1}+\Sigma_{2}}{2}$.
-The Bhattacharyya intuitively consists of two additive terms.
+The Bhattacharyya consists of two additive terms.
 The first term is the squared Mahalanobis distance (Mahalanobis, 1936).
@@
 The $\det\Sigma$ is the so-called generalized variance index (GVI) (Wilks, 1932), which provides a multivariate extension of the standard statistical variance measure based on the mean squared deviation.
-To calculate the BCD, we calculate both the means ($\mu$) and the variance-covariance matrix ($\Sigma$) of the embedding matrix $\mathbf{X}$ [^11] As the calculation of the dot product $XX^{\prime}$ is computationally unfeasible, we instead use a maximum likelihood estimator of the empirical covariance matrix using the procedure implemented in scikit learn.
-In tests on a subsample of the data, we confirmed that the approximate is very close to the analytical solution of the variance-covariance matrix.
+To calculate the BCD, we calculate both the means ($\mu$) and the variance-covariance matrix ($\Sigma$) of the embedding matrix $\mathbf{X}$.
+[^11] As the calculation of the dot product $XX^{\prime}$ is computationally unfeasible, we instead use a maximum likelihood estimator of the empirical covariance matrix using the procedure implemented in scikit learn.
+In tests on a subsample of the data, we confirmed that the approximation is very close to the analytical solution of the variance-covariance matrix.
 The diagonal elements of the variance-covariance matrix capture the dispersion of Tweets along the embedding dimensions.
@@
 Given that this equally affects all components of the BCD, this has no bearing on our results.
-In this way, BCD provides a unidimensional measure that summarizes the overall distortions of the embeddings matrix $\mathbf{X}$.
+In this way, BCD provides a unidimensional measure that summarizes the overall distortions of the embedding matrix $\mathbf{X}$.
 The BCD has several properties that make it a desirable measure for our application:
@@
 The BCD allows us to approximate the distortions of the embedding space in a computationally feasible manner, a crucial aspect given the expansive nature of online platforms and our data.
-[^13] Many other alternative methods, such as the ones that involve the computation of convex hulls, become computationally infeasible in high dimensions settings.
+[^13] Many other alternative methods, such as the ones that involve the computation of convex hulls, become computationally infeasible in high-dimensional settings.
+Advantages over Potential Alternative Metrics
+Readers familiar with the broader natural language processing literature may wonder why we propose a new measure to quantify content distortions, rather than relying on more commonly used approaches.
+The two most plausible candidates could be to calculate the cosine similarity of the embedding vectors or use topic models to analyse how the topic distribution has shifted.
+While these techniques are very useful and we will use a topic model later in our analyses, our approach has several advantages compared to these alternative measures.
+First, our measure is computationally significantly cheaper to construct from the embedding space than cosine-based measures, which is particularly relevant given the large size of social media platforms.
+In comparison, the required number of calculations and memory grow quadratically in the number of observations for the cosine similarity.
+Secondly, both cosine similarity and topic models may fail to detect semantic changes in online content, despite the social media content having undergone significant shifts.
+In Appendix C, we show that the average similarity of tweets remains virtually unchanged when toxic content is removed.
+Similarly, topic models do not account for the differences in opinions expressed within a topic (e.g., pro- and anti-immigration tweets).
+It would therefore be possible to remove one side from the political debate without necessarily shifting the topic composition.
+For example, if one political side accounted for half of each online topic, we could remove this side completely from each topic without altering the underlying distribution of topics.
+In contrast, our approach provides a holistic measure for the total extent of semantic shifts.
 3 Results
-In the following, we use the BCD to establish four sets of results.
-First, we analyze how the content is distorted when we remove highly toxic Tweets from the data.
+In the following, we use BCD to establish three sets of results.
+First, we analyze the extent to which online content is shifted when we remove highly toxic Tweets from the data.
 This analysis allows us to simulate the effect of more stringent content moderation.
-Second, we demonstrate that the content moderation-induced distortions are not driven by the toxic language of the Tweets.
-Third, we investigate what underlies the changes in the BCD.
-Fourth, we propose an alternative approach to content moderation that reduces toxicity while preserving the content of Tweets.
+Second, we benchmark the magnitude of these shifts and show that the content-moderation-induced shifts are comparable to the removal of entire topics from the online debate.
+Third, we demonstrate that content moderation-induced distortions can be reduced by rephrasing highly toxic Tweets in a way that removes the toxic language while maintaining the original content.
+The last finding suggests that the documented shifts in online content are not exclusively driven by toxic language, but rather stem from the removal of specific topics from the online debate.
 3.1 Removal of Toxic Content and Distortions of Online Discourse
-We simulate the effects of more stringent content moderation by removing toxic Tweets based on varying toxicity thresholds from our data.
-We then analyze if removing toxic Tweets leads to distortions of online content as measured by the BCD.
-We will compare these content-moderation-induced changes in the semantic space to changes in the BCD if we instead remove the same number of Tweets from the data at random.
-Note that this analysis intentionally abstracts from any form of behavioral reaction on the side of the users, as we are interested in the direct effects of content moderation.
-This analysis should provide a realistic content moderation benchmark as the Perspectives API is used in real-world applications, and several studies of content moderation have used toxicity thresholds to delineate toxic content (e.g., Gehman et al., 2020; Han and Tsvetkov, 2020; Rieder and Skop, 2021; Hede et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022).
-The results from this analysis are presented in Figure 1.
+In our initial analysis, we simulate the effects of more stringent content moderation by removing toxic Tweets based on varying toxicity thresholds from our data.
+This analysis provides a realistic content moderation benchmark as the Perspectives API is used in real-world applications, and several studies of content moderation have used toxicity thresholds to delineate toxic content (e.g., Gehman et al., 2020; Han and Tsvetkov, 2020; Rieder and Skop, 2021; Hede et al., 2021; Jiménez Durán, 2022; Beknazar-Yuzbashev et al., 2022).
+We then use BCD to analyze whether removing toxic Tweets leads to distortions of online content.
+We compare these content-moderation-induced changes to a baseline by removing the same number of Tweets from the data at random.
+Note that this analysis intentionally abstracts from any behavioral user reaction, as we are interested in the direct effects of content moderation.
+For example, the removal of toxic content could allow non-toxic users to more freely express their opinions, or toxic users could be deterred from using toxic language.
+In our analysis, we aim to isolate the first-round effects of content moderation; however, our measure would also be well-suited to characterize any behavioral reactions.
+The results from our first analysis are presented in Figure 1.
 The x-axis indicates the toxicity threshold above which we remove Tweets.
@@
 The y-axis shows our measure of the distortion of the semantic space after the removal of toxic content relative to the original semantic space as measured by the BCD.
-It is immediately apparent that the removal of toxic content distortions to the semantic space (orange line).
-The findings indicate that distortions become more severe the lower we set the toxicity threshold.
+It is immediately apparent that the removal of toxic content leads to distortions to the semantic space (orange line).
+Intuitively, the findings indicate that changes become more severe the lower we set the toxicity threshold.
 Importantly, removing Tweets from the data at random does not impact the BCD (blue line).
-This highlights that the increase in the BCD is not a mechanical consequence of a smaller sample of Tweets but rather driven by the changing composition of online discourse due to content moderation.
-We interpret this result as evidence that removing toxic distorts online content as measured by the embedding space.
+This highlights that the increase in the BCD is not a mechanical consequence of a smaller sample of Tweets but rather is driven by the changing composition of online discourse due to content moderation.
+This finding suggests that the removal of toxic online content leads to significant shifts in the embedding space.
 Figure 1: Content Distortions and Removal of Toxic Content
@@
 The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
-To benchmark the resulting distortions, we compare them to an approximation of the maximum possible distortion one could create by removing a specific number of Tweets.
-For this exercise, we remove the Tweets with the largest Euclidean distance from the centroid of the semantic space.
-Appendix Figure 1 shows the results from this exercise.
-We find that the removal of toxic content reaches around 20% of the maximum possible distortion.
-In Section 3.2, we provide additional benchmarks of the distortions relative to the direct removal of topics.
-Robustness
-We conduct several robustness exercises for this finding.
-First, we reproduce our finding using alternative embeddings based on the widely-used RoBerta and DeBerta models (see Figure C.2).
-Second, we base content moderation on the alternative Toxicity dimensions from the Perspectives API, the toxicity scores from the Detoxify classifiers (Hanu and Unitary team, 2020), and the Moderation API from OpenAI (see Figure C.3).
-Lastly, we repeat our analysis using samples of 1 million: 1) general Tweets (without filtering for political content), 2) German Tweets, and 3) Italian Tweets (see Figure C.4).
-We discuss all of these robustness checks in greater detail in Appendix C.
-To summarize, none of these changes make any qualitative difference to our findings.
-Independent of the embedding model, the toxicity scores, or the sample, removing toxic content reduces the plurality of online discourse.
-Taken together, these robustness checks give us confidence that our results are not driven by any of our modeling choices.
-3.2 What Drives Distortions to Online Content
-We next aim to shed light on the factors that drive the content-moderation-induced distortions of the semantic space we have documented.
-First, we investigate the individual components of the BCD and show that both the mean as well as the variance of the embedding space get distorted by content moderation.
-Second, we show that the increase in the BCD is not driven by the use of toxic language itself.
-Third, we provide further evidence for the content-moderation-induced distortions of the topics of online debate.
+Benchmarking the Magnitude of Distortions
+Next, we provide two benchmarks to better understand the magnitude of the content-moderation-induced alterations of the semantic space: 1) an upper bound for BCD, and 2) the effect of topic-based removal.
+First, we compare changes in the BCD when removing Tweets based on their toxicity scores versus removing an equal number of Tweets with the greatest Euclidean distance from the centroid of all embedded Tweets.
+Given the definition of Bhattacharyya distance, removing Tweets furthest from the distribution’s center approximates the upper bound of BCD changes possible by removing a specific number of Tweets.
+Figure 2 shows the changes in BCD when removing Tweets with toxicity scores above a certain threshold and compares this to removing an equal number of Tweets based on their distance from the centroid.
+The results indicate that removing toxic Tweets increases BCD by approximately 20% of the maximum possible increase from removing Tweets with a large Euclidean distance.
+Figure 2: Benchmarking BCD
+Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis.
+The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random.
+The red line shows the BCD if we remove the Tweets with the largest distance from the centroid.
+The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
+Second, we provide an additional benchmark of the distortions relative to the direct removal of topics.
+In Panel (a) of Figure 3, we analyze how much the semantic space changes if we directly delete Tweets discussing specific topics from the data.
+For this figure, we calculate the BCD for cases in which we randomly remove 1, 2, 3, 4, or 5 topics from the data.
+In each case, we implement 25 random draws of the indicated number of topics.
+We then remove Tweets from the selected topics from the data and calculate the BCD.
+The figure then reports the average BCD resulting from the 25 draws.
+We also report the average share of Tweets that remain in the data after the removal of topics in brackets on the x-axis.
+Intuitively, we observe that the BCD increases progressively as additional topics are removed.
+This exercise also allows us to provide another benchmark for the content-moderation-induced distortions we documented in Figure 1.
+We find that removing all Tweets with a toxicity score above 0.8 from the data is akin to removing four topics from the data at random.
+Both of these interventions result in a BCD of approximately 0.7.
+However, it is also worth highlighting that removing four topics deletes twice as many Tweets from the data (6%) as content moderation with a threshold of 0.8 (3%).
+This suggests that the per-Tweet distortions of the semantic space are more severe when content is removed based on toxicity.
+In Panel (b) of Figure 3, we additionally show the topics that are affected most by content moderation.
+Perhaps unsurprisingly, we find that Tweets containing insults against political figures are most affected by content moderation.
+[^14] Note that even highly uncivil insults against public figures represent a form of opinion expression.
+Moreover, we also find that Tweets discussing the topics of “Racism”, “Trump”, or the “Black Lives Matter (BLM)” movement are heavily affected by content moderation, thereby distorting the topic composition of online debate.
+[^15] The finding on the removal of discussions of racism aligns well with the work of Lee et al.
+(2024).
+It is also important to highlight that our results do not hinge on shifts in any one topic, i.e., the results are virtually identical if we do not consider Tweets containing insults (see Appendix Figure C.6).
+Furthermore, we can demonstrate that content moderation distorts the semantic space, even within specific topics (unreported).
+Figure 3: Content Moderation and Topic Shifts
+(a) BCD and Removal of Topics
+(b) Topics Shifts
+Notes: The figure visualizes the effect of content moderation on changes in topics as generated by the Top2Vec algorithm.
+Panel (a) shows that removing 1, 2, 3, 4, or 5 topics from the data at random leads to increases in the BCD.
+Panel (b) shows which topics are most heavily affected by content moderation.
+The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
 Decomposing the Bhatacharya Distance
+As the next step, we develop a better understanding of the statistical moments (mean and variance) that underlie increases in the BCD.
 As previously discussed, the BCD consists of two additive components (see eq.
 1), the first of which captures distortions to the mean, while the latter captures distortions to the variance.
-We analyze to what extent each of these components contributes to the previously documented increases in the BCD.
+We analyze the extent to which each of these components contributes to the previously documented increases in the BCD.
 The results from this analysis are shown in Panel (a) of Figure 4.
-We observe that while both the first and the second component contribute to the increase in the BCD, the shift in the variance explains around 96% of the increase in the BCD, while the shift in the mean only explains 4%.
+We observe that while both the first and second components contribute to the increase in the BCD, the shift in variance explains approximately 96% of the increase in the BCD, whereas the shift in the mean only explains 4%.
 Figure 4: Decompostion of Bhatacharya Distance
@@
 The results suggest that content moderation appears to remove outliers in the semantic space.
-Removing such “outlier” content is arguably costly for the plurality of online speech.
-Accounting for the Toxicity Dimension of the Embedding Space
-The content moderation-induced changes in the semantic space could be driven by two factors.
-On the one hand, the toxic language itself could occupy a particular part of the semantic space.
-Removing toxic content would then mechanically increase the BCD, but this would not necessarily indicate a “real” distortions to online content.
-On the other hand, toxic Tweets might discuss specific issues and topics that are underrepresented online, independently of their toxicity.
-We investigate this question by removing the toxicity component from the embedding space.
-To do so, we build on the literature on the debiasing of embeddings (e.g., Bolukbasi et al., 2016; Liang et al., 2020) and create orthogonal projections of the embeddings matrix $\mathbf{X}$ with respect to the Toxicity scores.
-Our baseline approach uses linear regressions of the following form:
-$x_{d}=\alpha+\mathbf{Tox^{\prime}\beta}+\epsilon_{d}$
-where $x_{d}\in\mathbf{X}$ is one of the $D$ embedding dimensions.
-$\mathbf{Tox}$ is a matrix containing 1000 indicators for the permilles of the toxicity distribution.
-[^14] We chose a non-parametric transformation of the toxicity scores to flexibly account for any potential non-linearities.
-As we show in a robustness check, the findings are almost identical if we instead residualize linearly with regard to the toxicity score.
-We estimate these regressions for all embedding dimensions $d\in D$ and replace the embedding dimensions with the regression residuals.
-This leaves us with a residualized matrix $\mathbf{\tilde{X}}$, which is orthogonal to the toxicity scores.
-In other words, we remove any variation in the embeddings of Tweets that can be explained by their toxicity.
-Any remaining variances in the embeddings should, therefore, capture differences in Tweets’ content independent of their toxicity.
-We then repeat our previous analysis based on the residualized embedding matrix $\mathbf{\tilde{X}}$.
-Figure 3 visualizes the two different approaches to account for a Tweet’s toxicity.
-More specifically, we use regressions and either residualize the embeddings using permilles or the linear toxicity score.
-We find that removing the toxicity component from the embedding space hardly makes any difference to our findings.
-The overall patterns are very close to our original results.
-The BCD sharply increases once toxic Tweets are removed from the data.
-This suggests that the observed changes in the BCD do not appear to be driven by the toxicity of the Tweets but rather by their content.
-Figure 3: Controlling for Toxicity
-Notes: The figure shows the BCD, derived from toxicity-debiased Tweet embeddings, after the exclusion of toxic and random Tweets from the sample.
-Linear and Percentile Residualization adjust each dimension of Tweets embeddings by using the residuals of the regression of the embedding values against the toxicity score and toxicity permilles of Tweets.
-Content-Moderation-Induced Topic Shifts
-As the last piece of evidence on the content-moderation-induced shifts in online content, we analyze how the topic composition of Tweets is shifting with increased content moderation.
-Additionally, we provide evidence that, in line with our original motivation, the removal of topics is associated with increases in the BCD.
-The latter result also allows us to benchmark the content-moderation-induced increases in the BCD relative to the direct removal of topics from the embedding space.
-In Panel (a) of Figure 4, we show the topics that would be affected most by content moderation.
-Perhaps unsurprisingly, we find that Tweets containing insults against political figures are most affected by content moderation.
-[^15] Note that even highly uncivil insults against public figures represent a form of opinion expression.
-Moreover, we also find that Tweets discussing the topics of “Racism”, “Trump”, or the “Black Lives Matter (BLM)” movement are heavily affected by content moderation, thereby distorting the topic composition of online debate.
-[^16] The finding on the removal of discussions of racism aligns well with the work of Lee et al.
-(2024).
-It is also important to highlight that our results do not hinge on shifts in any one topic, i.e., the results are very virtually identical if we do not consider Tweets containing insults (see Appendix Figure C.5).
-Further, we can show that content moderation distorts the semantic space even within topics (unreported).
-Figure 4: Content Moderation and Topic Shifts
-(a) Topics Shifts
-(b) BCD and Removal of Topics
-Notes: The figure visualizes the effect of content moderation on changes in topics as generated by the Top2Vec algorithm.
-Panel (a) shows which topics are most heavily affected by content moderation.
-Panel (b) shows that removing 1, 2, 3, 4, or 5 topics from the data at random leads to increases in the BCD.
-The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
-In Panel (b), we lastly show that removing topics leads to increasing changes in the BCD.
-For this figure, we calculate the BCD for the cases in which we remove 1, 2, 3, 4, or 5 topics from the data at random.
-For this exercise, we implement 25 random draws of the indicated number of topics.
-We then remove Tweets from the selected topics from the data and calculate the BCD.
-The figure then reports the average BCD resulting from the 25 draws.
-We also report the average share of Tweets that remain in the data after the removal of topics in brackets on the x-axis.
-We find that the BCD increases progressively with the removal of additional topics.
-This exercise also allows us to benchmark the content-moderation-induced distortions we documented in Figure 1.
-We find that removing all Tweets with a toxicity score above 0.8 from the data is akin to removing four topics from the data at random.
-Both of these interventions result in a BCD of ca.
-0.7.
-It is also worth highlighting that removing four deletes twice as many Tweets from the data (6%) as content moderation with a threshold of 0.8 (3%).
-This suggests that the per-Tweet distortions of the semantic space are more severe when removing content based on toxicity.
-Taken stock, the previous exercises suggest that the reduction in BCD are not driven by the toxic language itself, but rather the fact that removing Tweets with toxic language as a byproduct 1) shifts the mean of the embeddings space, 2) reduces the variance of the semantic space, and 3) removes topics from the online debate.
-In the last part of the paper, we propose an alternative approach to mitigate some of the content-moderation-induced distortions.
-3.3 An Alternative Approach to Content Moderation
-We propose an alternative approach to tackle the issue of online toxicity that has the potential to mitigate some of the distortions in online discourse.
-Instead of removing toxic content outright, content moderators could instead use the language generation capabilities of large language models and rephrase the message of Tweets using less toxic language.
-This makes it possible to reduce the toxicity of online content while keeping the original content as far as possible intact.
+Removing such “outlier” content is arguably costly for the plurality of online speech as they represent positions that are expressed less frequently online.
+Robustness
+We conduct several robustness checks to verify our findings.
+First, we reproduce our finding using alternative embeddings based on the widely used RoBERTa, DeBERTa, and DistilBert-Base-Multilingual-Cased-V2 models (see Figure C.2).
+Second, we base content moderation on the alternative Toxicity dimensions from the Perspectives API, the toxicity scores from the Detoxify classifiers (Hanu and Unitary team, 2020), and the Moderation API from OpenAI (see Figure C.3).
+Third, we show that our results are very similar when we weight Tweets by their user engagement, as proxied by the number of Retweets.
+Lastly, we repeat our analysis using samples of 1 million: 1) general Tweets (without filtering for political content), 2) German Tweets, and 3) Italian Tweets (see Figure C.5).
+We discuss all of these robustness checks in greater detail in Appendix C.
+To summarize, none of these changes makes any qualitative difference to our findings.
+Independent of the embedding model, the toxicity scores, or the sample, removing toxic content reduces the plurality of online discourse.
+Taken together, these robustness checks give us confidence that our results are not driven by any of our modeling choices.
+3.2 What Drives Distortions to Online Content
+The previously documented distortions may arise for two reasons.
+On the one hand, they could reflect a mechanical shift in the semantic space caused by the removal of toxic language.
+Abstracting from debates over what constitutes toxicity, such shifts would arguably not be costly, since they merely reflect the removal of content already deemed unacceptable for the platform.
+On the other hand, toxic Tweets might discuss specific issues and topics that are underrepresented online using inflammatory language.
+In this case, removing such Tweets would distort the online debate beyond the elimination of the toxic language itself.
+We evaluate these competing hypotheses using two complementary approaches.
+First, we use large language models (LLMs) to rephrase Tweets, stripping away toxic language while preserving the underlying message.
+We show that replacing Tweets with their rephrased counterparts (instead of removing them entirely) mitigates some of the distortions to the semantic space.
+This result highlights that the distortions of the semantic space are not entirely driven by the removal of toxic language.
+Furthermore, this exercise demonstrates the potential of our measure to benchmark different content moderation strategies against one another.
+Second, building on the literature on debiasing text embeddings (e.g., Bolukbasi et al., 2016), we demonstrate that moderation-induced distortions are not attributable to a specific spatial positioning of toxic language in the embedding space, but rather to changes in online content.
+Specifically, we construct projections of the embedding space orthogonal to toxicity scores (i.e., embeddings that assign the same vector representation to Tweets regardless of their toxicity).
+We find that content moderation still induces distortions in this orthogonalized embedding space, suggesting that the distortions are not merely an artifact of removing toxic language.
+Furthermore, we show that removing toxic Tweets in the orthogonalized embedding space, as opposed to removing their rephrased counterparts in the original embedding space, produces nearly identical BCDs.
+Rephrasing Tweets
+In the first analysis, we show that it is possible to reduce the toxicity of online content while creating smaller distortions to the semantic space.
+We thereby highlight that there is content that can be salvaged from toxic Tweets.
+For this analysis, we propose an alternative approach to address the issue of online toxicity, which has the potential to mitigate some of the distortions in online discourse.
+Instead of removing toxic content outright, content moderators could use the language generation capabilities of LLMs to rephrase the message of Tweets using less toxic language.
+This transformation reduces the toxicity of online content while preserving the original content as much as possible.
 [^16] The detoxification of online content has recently also become a task tackled by computer scientists in the TextDetox 2024 competition.
@@
 There might well be content and thoughts that should not be admitted to be shared on online platforms.
-Our argument rather is that there exist significant amounts of content that violate the content standards of social media platforms (e.g., grave insults) but, at the same time, might contain political expression or information that might be worth preserving.
+Some posts are beyond salvaging.
+Our argument, rather, is that a substantial share of content violating platform standards (e.g., grave insults) or toxicity thresholds may nonetheless carry political expression or information worth preserving.
 In such cases, rephrasing might be a suitable approach.
@@
 We provide additional details on the rephrasing prompt in Section B.2.
-We also show examples of rephrased Tweets in Table A.3.
-Overall, GPT4o-mini can rephrase Tweets in a significantly more civil manner while keeping the main message intact.
-Rephrasing reduces the toxicity of Tweets from 0.71 to 0.26.
+We also show examples of rephrased Tweets in Table 1.
+Overall, GPT4o-mini can rephrase Tweets in a significantly more civil manner while retaining the main message.
+Rephrasing reduces the toxicity of Tweets from 0.71 to 0.26, while maintaining very similar content (i.e., the average cosine similarity between rephrased Tweets and the originals is 0.97).
 We then compare the BCD for two content moderation strategies.
@@
 The results from this analysis are shown in Figure 5.
-We find that in contrast to removal, the rephrasing of Tweets leads to smaller changes in the BCD.
-The fact that the gap is widening with lower content-moderation thresholds strikes us as intuitive as rephrasing highly toxic content that only contains insults leads to similar distortions than outright removal.
-Only once we consider content that contains other content independent of the toxic language, rephrasing achieves its intended goal of preserving the original content.
-This suggests that such an approach can reduce the toxicity of online content with fewer distortions to online content.
+We find that, in contrast to removal, the rephrasing of Tweets leads to smaller changes in the BCD.
+If the content changes were solely driven by the presence of toxic language, replacing Tweets with a rephrased version should cause the same change to the semantic space.
+However, since we find that rephrasing can mitigate some of the distortions, it suggests that the observed distortions are not merely artifacts of toxic language itself.
+The fact that the gap is widening with lower content-moderation thresholds strikes us as intuitive, as rephrasing highly toxic content that only contains insults leads to similar distortions as outright removal.
+These are the posts that cannot be salvaged.
+Only once we consider Tweets that contain other content, independent of the toxic language, does rephrasing achieve its intended goal of preserving the original content.
+This suggests that rephrasing can reduce the toxicity of online content with fewer distortions.
+Note that we would not expect the BCD of the rephrased Tweets to be zero, as rephrasing introduces changes in the language of Tweets.
 Figure 5: Content Plurality and Rephrasing of Toxic Content
@@
 The blue line shows the BCD if toxic Tweets are rephrased.
+Accounting for the Toxicity Dimension of the Embedding Space
+As a second analysis, we directly remove the toxicity component from the embedding space.
+To do so, we build on the literature on the debiasing of embeddings (e.g., Bolukbasi et al., 2016; Liang et al., 2020) and create orthogonal projections of the embeddings matrix $\mathbf{X}$ with respect to the Toxicity scores.
+In other words, we remove any variation in the embeddings of Tweets that can be explained by their toxicity.
+Any remaining variances in the embeddings should, therefore, capture differences in the content of the Tweets, independent of their toxicity.
+Initially, we construct a “toxicity dimension” by calculating the average vector differences between the embeddings of highly toxic Tweets and their rephrased counterparts.
+Next, we use Principal Component Analysis (PCA) to identify the 10 principal vectors that characterize toxicity.
+The final step involves projecting the embedding space so that it is orthogonal with respect to the toxicity subspace defined by the principal components.
+This procedure produces an orthogonalized matrix $\mathbf{\tilde{X}}$, which is uncorrelated with the toxicity scores.
+In Online Appendix C, we also present an alternative method to remove the toxicity dimension from the embedding space based on regression residuals, which yields similar results.
+We then repeat the previous analysis using the orthogonalized embedding matrix $\mathbf{\tilde{X}}$.
+Figure 6 compares the BCD obtained when removing Tweets from the original embedding space (orange line) and the orthogonalized embedding space (blue line).
+We also report the BCD for the removal of rephrased Tweets (red line).
+[^17] In this exercise, all Tweets are first replaced with their rephrased versions to define the baseline sample.
+Afterwards, Tweets are removed based on their pre-rephrasing toxicity scores.
+Three key findings emerge from this analysis.
+First, even in the orthogonalized embedding space, we continue to observe substantial, albeit somewhat smaller, changes in the BCD.
+This again suggests that the observed shifts in BCD arise to a considerable extent from the underlying content of Tweets rather than their toxic language.
+Second, the BCD resulting from the removal of orthogonalized Tweets is virtually identical to that obtained by removing their rephrased counterparts.
+This demonstrates that our orthogonalization procedure effectively removed the toxicity dimension, bringing the embeddings close to those of the rephrased Tweets.
+Third, removing rephrased Tweets produces more pronounced changes to the embedding space than replacing toxic Tweets with their rephrased versions (see Figure 5).
+Intuitively, this underscores that removing content introduces larger distortions to the embedding space than rephrasing.
+Figure 6: Controlling for Toxicity
+Notes: The figure shows the BCD, derived from the original embedding space (orange line) and the orthogonalized embedding space (blue line).
+We also report the BCD for the removal of rephrased Tweets (red line).
 4 Conclusion
-This paper proposes and validates a new methodology to measure content-moderation-induced distortions to online content.
-This new methodology enables us to, for the first time, document the impact of the removal of toxic content on online discourse.
+This paper proposes and validates a new methodology for measuring content-moderation-induced distortions in online content.
+This new methodology enables us, for the first time, to document the impact of removing toxic content on online discourse.
 Given the crucial importance and heated nature of the debate surrounding this issue, it is important to be clear about what our measure achieves and what it does not capture.
-The measure captures the distortions in the embedding space of the content that is circulating on online platforms.
+The BCD provides a heuristic measure for changes in the embedding space of content circulating on online platforms.
 As the embeddings measure many semantic characteristics of a text, changes in the mean and the variance of the embeddings provide insight into the extent of the semantic distortions.
-That being said, our measure does not, nor is it the goal of this paper, provide a universal measure of online content and free speech.
-Given the complex and multifaceted nature of the concept of free speech that encompasses different philosophical and legal standpoints (see for example Warburton, 2009, for a review), no one measure could ever capture all facets of online discourse.
-In the same way, despite extensive efforts by the research community, no measure of hate speech could ever capture the importance of cultural context and changing societal norms (e.g., Brown, 2017).
-Nonetheless, automated hate speech detection tools have proven useful to rain in toxic speech and are hence widely deployed online.
-As online platforms and regulators inevitably face trade-offs when it comes to the moderation of online content, we believe it is important to have measures that are also able to capture the cost of content moderation.
-We believe that by shedding light on the trade-offs involved in content moderation, our measure provides a fundamental advancement in the application of NLP tools for the problem of content moderation and highlights a highly fruitful direction for future research.
+That being said, our measure does not provide a universal measure of online content and free speech (nor is it the goal of this paper).
+Given the complex and multifaceted nature of the concept of free speech, which encompasses different philosophical and legal standpoints (see, for example, Warburton, 2009, for a review), no single measure can ever capture all facets of online discourse.
+Similarly, despite extensive efforts by the research community, no measure of hate speech can ever fully capture the importance of cultural context and changing societal norms (e.g., Brown, 2017).
+Nonetheless, automated hate speech detection tools have proven helpful to rein in toxic speech and are hence widely deployed online.
+As online platforms and regulators inevitably face trade-offs when it comes to moderating online content, we believe it is essential to have measures that also capture the cost of content moderation.
+We believe that by shedding light on the trade-offs involved in content moderation, our measure represents a fundamental advancement in the application of NLP tools for content moderation and highlights a highly fruitful direction for future research.
 Appendix A Additional Details on Data
@@
 This particular sample selection bears two pivotal advantages.
-Firstly, it affords us the opportunity to construct a comprehensive and representative overview of Twitter activity across the United States.
+Firstly, it allows us to construct a comprehensive and representative overview of Twitter activity across the United States.
 Secondly, these users have been actively engaged on Twitter for several years, and we do not face problems with composition changes.
-Consequently, we were able to conduct an insightful analysis based on a good approximation of the content that circulated on Twitter.
-The initial step in our data collection involved retrieving the Tweets posted by each user within the (Siegel et al., 2021) sample.
+Consequently, we were able to conduct our analysis based on a good approximation of the content that circulated on Twitter.
+The initial step in our data collection involved retrieving the Tweets posted by each user within the Siegel et al.
+(2021) sample.
 In totality, our dataset encompassed 399 Million Tweets spanning from 2014 to the commencement of 2022, along with corresponding user profile information.
@@
 Each subset included 200 shared Tweets to evaluate inter-coder reliability.
-We evaluated agreement using Cohen’s kappa coefficient, which ranges from 0 to 1, where 1 indicates perfect agreement and values near 0 suggest agreement no better than chance.
+We evaluated agreement using Cohen’s kappa coefficient, which ranges from 0 to 1, where 1 indicates perfect agreement and values near zero suggest agreement no better than chance.
 The average Cohen’s kappa score across all coder pairs was 0.76.
@@
 This state-of-the-art API assigns a toxicity score ranging from 0 to 1 across six distinct dimensions: general toxicity, severe toxicity, identity attacks, insults, profanity, and threats.
-The scores can be approximately interpreted as the probability of a randomly chosen user classifying content as toxic.
+The scores can be approximately interpreted as the probability that a randomly chosen user will classify content as toxic.
 For example, a score of 0.8 means that around 80% of users would judge the content to be toxic.
 The Perspective API performs well in classifying toxic text and can assess Tweets in several languages, including English, Spanish, French, German, Portuguese, Italian, and Russian.
-In our dataset, English, unsurprisingly, is the predominant language, and we restrict our analysis to English-speaking Tweets.
+In our dataset, English is, unsurprisingly, the predominant language, and we restrict our analysis to English-speaking Tweets.
 Instances where the API did not assign toxicity scores were primarily due to the absence of textual content, such as Tweets containing only hyperlinks.
-We provide examples of highly toxic Tweets in Table A.3.
+We provide examples of highly toxic Tweets in Table 1.
 The examples indicate that the Perspective API accurately identifies toxic content.
@@
 Data points are grouped into 20 bins of equal size.
-Table A.3: Examples Toxic Tweets
-Tweet Text Toxicity Score Rephrased Tweet Again, to the 60+ million so-called Americans who voted for this racist hateful bullshit: FUCK YOU.
-0.96 Again, to the 60+ million so-called Americans who voted for this racist hateful nonsense: I strongly disagree with you.
-@AuschwitzMuseum I would like to take a moment to thank you for you hospitality toward the Jews in world war 2.
-Fuck Jews 9/11 #hitlerbaby!
-0.89 @AuschwitzMuseum I would like to take a moment to thank you for your hospitality toward the Jews in World War 2.
-I have strong feelings about this topic.
-#hitlerbaby!
-Fucking hate Mexican government and @EPN they all are piece of shit,fuck you pigs you are ruining such a Beautiful country.
-0.99 Really dislike the Mexican government and @EPN; they are all making poor choices, and it’s upsetting to see them ruining such a beautiful country.
-The Muslim Prophet Muhammad was a pedophile, rapist, murderer.
-#Cult #Islam #TrumpOnCNN 0.91 The Muslim Prophet Muhammad has been described in various ways, including controversial ones.
-#Cult #Islam #TrumpOnCNN @michellemalkin Put all Obamas people on an island together gays,blacks,Mexicans and uneducated birth control women Wow !
-What an island 0.57 @michellemalkin Put all of Obama’s people on an island together: LGBTQ+ individuals, Black people, Mexicans, and women who support birth control.
-Wow!
-What an island!
-Notes: This table presents examples of toxic Tweets, the toxicity scores from the Perspectives API, and the rephrased non-toxic versions.
 Additional Details on Topic Model
-To understand the topic composition of our underlying Twitter data and the changes that occur with increased content moderation, we train a Top2Vec topic model (Angelov, 2020) on our sample of political Tweets.
-Top2Vec combines embeddings from pre-trained transformer models with unsupervised clustering algorithms to derive highly interpretable topics.
-By relying on pre-trained embeddings, Top2Vec can handle large and diverse datasets more effectively than traditional topic models like Latent Dirichlet Allocation (LDA) (Blei et al., 2003) and provide a more nuanced description of the underlying topics, especially when dealing with short texts like Tweets.
-A further advantage of Top2Vec is that it automatically chooses the number of clusters, which is a key challenge when training traditional topic models.
-The Top2Vec algorithm proceeds in three steps.
-First, the texts are transformed into embedding vectors based on pre-trained models.
-In our particular case, we make use of the universal-sentence-encoder model (Cer et al., 2018).
-Second, the embeddings are projected into lower dimensional space using UMAP (McInnes et al., 2018).
-This helps to overcome the sparsity of the higher dimensional space.
-Third, the reduced embeddings are clustered into topics using HDBSCAN (McInnes et al., 2017), which identifies the centroids for each topic vector.
-We specify that HDBSCAN should only create clusters with at least 1500 observations.
-Top2Vec creates a unique assignment between texts and topics.
-The topic words are then derived based on the word vectors that are closest to the topic centroid.
-We report the most important topics words as well as well as the assigned topic labels in the table below.
 Table A.3: Most Relevant Words by Topic
@@
 B.1. Additional Details on Embeddings
-To distill information from the raw text of Tweets into quantifiable data, we use three pre-trained models: BERTweet (Nguyen et al., 2020), RoBERTa (Liu et al., 2019), and DeBERTa (He et al., 2021).
-These models all built on the BERT-style Transformer architecture, which has become a staple in text classification and natural language understanding tasks.
+To distill information from the raw text of Tweets into quantifiable data, we use three pre-trained models: BERTweet (Nguyen et al., 2020), RoBERTa (Liu et al., 2019), DeBERTa (He et al., 2021), and DistilBert-Base-Multilingual-Cased-V2 (Reimers and Gurevych, 2019).
+These models all use on the BERT-style Transformer architecture, which has become a staple in text classification and natural language understanding tasks.
 The strength of these models comes from their utilization of an ”attention” mechanism, allowing them to evaluate words in the context of their surrounding text, thereby capturing the subtleties of language that are often lost in traditional analysis.
 We transform the raw Tweets into analyzable embeddings by tokenizing the text into its constituent word tokens.
-These tokens then serve as input to our models which produce numerical representations—or embeddings—of each Tweet.
+These tokens then serve as input to our models, which produce numerical representations—or embeddings—of each Tweet.
 These embeddings convey the semantic and syntactic nuances of the language used by Twitter users and form the backbone of our computational analysis.
@@
 This allows for a better interpretation of language nuances.
-Furthermore, DeBERTa’s enhanced mask decoder predicts masked tokens by utilizing their absolute positions, an improvement that aids in the pre-training process.
+Furthermore, DeBERTa’s enhanced mask decoder predicts masked tokens by using their absolute positions, an improvement that aids in the pre-training process.
 With these advancements, DeBERTa improved the model’s pre-training efficiency and improved the performance across a variety of downstream tasks, like the generation of new text that mimics human speech.
+DistilBert-Base-Multilingual-Cased-V2
+DistilBERT-Base-Multilingual-Cased-V2 (Reimers and Gurevych, 2019) is a lighter and faster variant of BERT, specifically trained to handle multiple languages in a single model.
+DistilBERT applies knowledge distillation during training.
+This process reduces the model’s size and increases inference speed.
+The multilingual-cased version was trained on multiple languages using cased text, ensuring that it preserves distinctions in capitalization, which can be semantically relevant in many languages.
+As a result, this model is particularly well-suited for cross-lingual applications.
 B.2. Additional Details on the Rephrasing of Tweets
@@
 The following section describes additional results and robustness checks.
-Specifically, we benchmark the content-moderation-induced distortions relative to the maximum possible distortion that can be created when removing a specific number of Tweets.
-Further, we repeated our main analysis using alternative 1) embedding models, 2) toxicity scores, and 3) samples.
-We also show that our results are not driven by Tweets belonging to the “Insult” topic as created by the Top2Vec topic model.
-Benchmarking BCD
-To benchmark the magnitude of the content-moderation-induced distortions, we compare changes in the BCD when removing Tweets based on their toxicity scores versus removing an equal number of Tweets with the greatest Euclidean distance from the centroid of all embedded Tweets.
-Given the definition of Bhattacharyya distance, removing Tweets furthest from the distribution’s center approximates the upper bound of BCD changes possible by removing a specific number of Tweets.
-Figure Figure C.1 shows the changes in BCD when removing Tweets with toxicity scores above a certain threshold and compares this to removing an equal number of Tweets based on their distance from the centroid.
-The results indicate that removing toxic Tweets increases BCD by approximately 20% of the maximum increase from removing Tweets with a large Euclidean distance.
-Figure C.1: Benchmarking BCD
-Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis.
-The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random.
-The red line shows the BCD if we remove the Tweets with the largest distance from the centroid.
-The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
+We begin by showing that the average similarity of Tweets is not affected by content moderation.
+Further, we repeated our main analysis using alternative 1) embedding models, 2) toxicity scores, 3) engagement weighting, and 4) samples.
+We also demonstrate that our results are not driven by Tweets belonging to the “Insult” topic as created by the Top2Vec topic model.
+Lastly, we provide evidence from an alternative method to account for the toxicity dimension of the embedding space.
+Removal of Toxic Content and Cosine Similarity
+In Figure C.1, we report the average cosine similarity of Tweets after sequentially removing those above varying toxicity thresholds.
+Given the very high computational requirements of the similarity calculation, we conduct this analysis for a random subset of 500,000 Tweets.
+The results show that average similarity remains largely unaffected by the exclusion of toxic content.
+In unreported analyses, we confirm that this pattern holds when we focus on a subset of Tweets with the highest and lowest pairwise similarity.
+Figure C.1: Removal of Toxic Content and Average Cosine Similarity
+Notes: The figure shows the average cosine similarity of Tweets after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis.
 Alternative Embeddings
-First, we reproduce our findings using the alternative embeddings described in Section B.1.
+As a first robustness check, we reproduce our findings using the alternative embeddings described in Section B.1.
 This test rules out that our findings are driven by the particularities of the specific transformer model we have chosen, even though BERTweet is one of the standard choices for the analysis of English-speaking Twitter data.
-For this robustness exercise, we created new embeddings based on the RoBERTa and DeBERTa models and reconstructed the BCD based on these embeddings.
+For this robustness exercise, we created new embeddings based on the RoBERTa, DeBERTa, and DistilBert-Base-Multilingual-Cased-V2 models and reconstructed the BCD based on these embeddings.
 The findings in Figure C.2 highlight that the findings are remarkably similar not only with regards to the overall patterns but also the magnitudes of the content-moderation-induced increases in the BCD.
 Figure C.2: Robustness: Alternative Embeddings
-Notes: The figure shows the BCD computed using embeddings generated by DeBerta and RoBerta, after the exclusion of toxic and random Tweets from the data.
+Notes: The figure shows the BCD computed using embeddings generated by DeBerta and RoBerta, and DistilBert-Base-Multilingual-Cased-V2 after the exclusion of toxic and random Tweets from the data.
 Alternative Toxicity Measures
@@
 Conversely, Panel (b) shows this measure after the removal of toxic Tweets, using toxicity scores generated by Detoxify and the OpenAI Moderation API.
+Engagement Weigting of Content
+As a third robustness test, we weight the Tweets in our data by their engagement as measured by the number of Retweets when calculating the BCD.
+This allows us to arguably better account for for the frequency with which users would encounter the toxic tweets.
+The results from this analysis are shown in Figure C.4.
+We find that weighting Tweets by their engagement has no bearing on our results and leads to very similar magnitudes for the BCD.
+Figure C.4: Robustness: Engagement Weighting
+Notes: The figure shows the BCD after excluding Tweets with a toxicity score exceeding the threshold shown on the x-axis.
+We weight Tweets by the number of retweets when calculating the BCD.
+The blue line illustrates the BCD when an equivalent number of Tweets is excluded from the dataset at random.
 Alternative Samples
@@
 Second, we use 1 million randomly drawn Tweets from a sample of politically interested German Twitter users.
-Third, we use 1 million randomly drawn Tweets from a sample of politically-interested Italian Twitter users.
+Third, we use 1 million randomly drawn Tweets from a sample of politically interested Italian Twitter users.
 To create embeddings for the German and Italian samples, we use the “bert-base-german-uncased” and “twitter-xlm-roberta-base” models, respectively.
@@
 The figure below reproduces Figure 1 when we remove all Tweets belonging to the “Insult” topic in advance.
-The results are virtually identical, suggesting again that content moderation also introduces distortions of the semantic space over and above purely removing uncivil content.
+The results are virtually identical, suggesting that content moderation also introduces distortions of the semantic space beyond simply removing uncivil content.
 Figure C.6: Content Distortions and Removal of Toxic Content (No Insults Topic)
@@
 The percentages in parentheses on the x-axis represent the proportion of Tweets retained relative to the original sample size.
+Alternative Method to Account for the Toxicity Dimension of the Embeddings
+Our alternative approach for the orthogonalization of the embeddings with respect to the toxicity scores uses linear regressions of the following form:
+$x_{d}=\alpha+\mathbf{Tox^{\prime}\beta}+\epsilon_{d}$
+where $x_{d}\in\mathbf{X}$ is one of the $D$ embedding dimensions.
+$\mathbf{Tox}$ is a matrix containing 1000 indicators for the permilles of the toxicity distribution.
+[^19] We chose a non-parametric transformation of the toxicity scores to flexibly account for any potential non-linearities.
+As we show in a robustness check, the findings are almost identical if we instead residualize linearly with regard to the toxicity score.
+We estimate these regressions for all embedding dimensions $d\in D$ and replace the embedding dimensions with the regression residuals.
+The resulting residualized matrix $\mathbf{\tilde{X}}$ is orthogonal to the toxicity scores.
+We then repeat our previous analysis based on the residualized embedding matrix $\mathbf{\tilde{X}}$.
+Figure C.7 visualizes the two different approaches to account for a Tweet’s toxicity.
+More specifically, we use regressions and either residualize the embeddings using permilles or the linear toxicity score.
+We again find that removing the toxicity component from the embedding space has little impact on our findings.
+The overall patterns are close to our original results.
+The BCD sharply increases once toxic Tweets are removed from the data.
+Figure C.7: Controlling for Toxicity
+Notes: The figure shows the BCD, derived from toxicity-debiased Tweet embeddings, after the exclusion of toxic and random Tweets from the sample.
+Linear and Percentile Residualization adjust each dimension of Tweets embeddings by using the residuals of the regression of the embedding values against the toxicity score and toxicity permilles of Tweets.